### v1

# Exploiting Domain-Specific Parallel Data on Multilingual Language Models for Low-resource Language Translation

Abstract
Neural Machine Translation (NMT) systems built on multilingual sequence-to-sequence Language Models (msLMs) fail to deliver expected results when the amount of parallel data for a language, as well as the language’s representation in the model are limited. This restricts the capabilities of domain-specific NMT systems for low-resource languages (LRLs). As a solution, parallel data from auxiliary domains can be used either to fine-tune or to further pre-train the msLM. We present an evaluation of the effectiveness of these two techniques in the context of domain-specific LRL-NMT. We also explore the impact of domain divergence on NMT model performance. We recommend several strategies for utilizing auxiliary parallel data in building domain-specific NMT models for LRLs.
keywords:
Neural Machine Translation , Domain divergence , multilingual Language Models , domain adaptation[label1]organization=School of Mathematical and Computational Sciences,addressline=Massey University, city=Albany, postcode=0632, state=Auckland, country=New Zealand
1 Introduction
The need to rapidly develop translation systems to disseminate information during pandemics, natural disasters, and other major events justifies the need to build domain-specific Neural Machine Translation (NMT) systems with any and all available parallel data. This is particularly true for low-resource languages (LRLs) that do not possess large amounts of parallel data for any single domain [1]. Therefore, when building a domain-specific NMT model for a given LRL pair, we face two possible scenarios: the considered (target, aka test) domain does not have any parallel data to be used in training, or it has a little amount of parallel data. In both cases, it is possible to make use of parallel data available for the same language pair, but in different (auxiliary) domains. This is not an unrealistic assumption - according to Tiedemann [2], hundreds of languages have parallel data from more than one domain.
Multilingual sequence-to-sequence Language Models (msLMs) such as mBART [3] have been shown to outperform vanilla Transformer models for the LRL-NMT task [4, 5, 6]. There are two possible ways to exploit auxiliary domain parallel data in building domain-specific LRL-NMT systems on top of msLMs: to further pre-train the msLM with self-supervised training objectives (continuous pre-training) [7] or to fine-tune the msLM with the NMT objective.
As shown in Figure 1, there are different ways to use auxiliary domain data in fine-tuning. These can be categorized as single-stage fine-tuning (FT) and Intermediate Task Transfer Learning (ITTL). Similarly, we can introduce another categorization based on the data availability of the target domain. In-domain refers to the case whether target domain has some parallel data for model training. Out-domain refers to the case where there is no such parallel data for the target domain. Note that out-domain testing scenario can be considered as zero-shot testing with respect to the target domain.
Variations of single-stage FT are: 1) single domain (vanilla) FT, where the msLM is fine-tuned once with data from one domain, and 2) multi-domain FT, where the msLM is fine-tuned once with data from multiple domains. In vanilla FT, if the target domain has parallel data, this data is used for fine-tuning (in-domain case). In multi-domain FT, when the target domain has parallel data (in-domain case), it is combined with auxiliary domain data. In ITTL, the msLM is fine-tuned with an intermediate task, before fine-tuning with the final task. It also has two variations: single-domain ITTL, where the intermediate task has data from one auxiliary domain, and multi-domain ITTL, where the intermediate task has data from multiple domains. In both cases, the final task (i.e. stage111We use the term stage interchangeably with task, i.e. final task or final stage, and intermediate task or intermediate stage.) fine-tuning is done using data from a single domain. If target domain has parallel data (in-domain case), the final task always uses this data. In the in-domain case, target domain data is also combined with auxiliary domain data in intermediate stage of multi-domain ITTL.
Currently, there is no comparative study on the impact of continuous pre-training and fine-tuning techniques when auxiliary domain data is used. Determining the method that best exploits auxiliary domain parallel data is beneficial for LRL-NMT researchers who are unable to run extensive experiments due to limited computational resources. Lee et al. [6] and Adelani et al. [8] identified that domain divergence in parallel datasets affects NMT performance. However, they did not conduct a detailed investigation into the impact of this factor. Khiu et al. [9] and our previous work [10] evaluated the impact of domain divergence only in the case of single domain ITTL. Given the above shortcomings, this paper intends to answer the following questions related to domain-specific LRL-NMT systems built on top of msLMs:
1.
Should we use auxiliary domain parallel data for fine-tuning, continuous pre-training, or for both?
2.
What is the impact of domain divergence between different parallel data sets on different FT setups?
In order to answer these questions, we conducted an extensive set of experiments with the aforementioned FT techniques as well as the bitext denoising pre-training technique [11]. We considered the LRLs Sinhala (si), Tamil (ta), Kannada (ka), and Gujarati (gu) as well as Hindi (hi) for English-centric NMT. We report results for the case where the target domain has some parallel data for training (in-domain) and where the target domain has no parallel data (out-domain). We quantify the impact of domain divergence using Jenson-Shanon Divergence (JSD).
Based on our analysis, we make the following observations in using parallel data from auxiliary domains when building NMT systems on top of msLMs for LRL pairs:
1.
When the parallel data set size is small (less than 50k, in our experiments), pre-training with bitext denoising yields no gains for both in- and out-domain setups.
2.
Multi-domain ITTL is the best for in-domain, but this utility diminishes when target dataset size increases. For out-domain, the best strategy depends on the domain relatedness and the size of parallel corpora.
3.
For both in- and out-domain setups, simply combining data from multiple domains does not help in ITTL, mainly due to the divergence between the domains they belong to.
2 Related Work
2.1 ITTL on msLMs for NMT
ITTL has roots in Transfer Learning, which has been extensively explored for NMT with vanilla Transformer models. However, as mentioned earlier, for LRLs, NMT systems built on top of msLMs outperform these traditional NMT techniques. Thus we do not discuss these techniques.
ITTL [12, 13, 14] refers to fine-tuning a pre-trained language model with the intermediate task(s), before fine-tuning on the final task. It has been extensively experimented with encoder-based LLMs such as BERT [12, 15], and has shown promising results even for LRLs [16].
With respect to ITTL for NMT, the closest to our work is Adelani et al. [8]. Similar to us, they explored single-domain ITTL and multi-domain ITTL, as well as multi-domain FT. However, they only considered the scenario where the auxiliary task has a large amount of data, while the target task has only a small amount of data. Moreover, they only used religious and news data and tested only for the in-domain setup. They did not quantify the impact of domain divergence. Our prior work [10] and Khiu et al. [9] considered only single-domain ITTL. However, they considered both in- and out-domain test setups and investigated the impact of domain divergence. Verma et al. [17] experimented with single-domain ITTL and multilingual ITTL (data from the same domain but in multiple languages). They found that the former is better, but did not quantify the impact of domain divergence.
2.2 Pre-training msLMs with Parallel Data
Reid and Artetxe [7] augmented the existing denoising objective in mBART with three new objectives: replace words in the noised sequence with a bilingual dictionary, predict the reference translation instead of the input sequence (bitext denoising), and a combination of the two. Kale et al. [18] used four denoising objective functions on mT5: translation language modeling, standard NMT, denoised NMT and denoised NMT with language model (LM). Chi et al. [19] used three cross-lingual objective functions on mT5: machine translation, translation pair span corruption, and translation span corruption. However, the latter two research did not consider the NMT task.
3 Methodology
The baseline for all our experiments is fine-tuning the msLM once with data from one domain (vanilla FT or single-domain FT). In the in-domain setup, this dataset comes from the target domain.
For continuous pre-training, we experimented with two of the Reid and Artetxe’s pre-training objective functions: bitext de-noising with auxiliary domain parallel data with and without monolingual denoising [11]. For monolingual denoising, we used the source and target side of the parallel data separately and did not use additional monolingual data as done by Reid and Artetxe [11], because our objective is only to exploit the use of parallel data. We also excluded their dictionary denoising, as many LRLs still do not have sufficient bilingual dictionaries [20]. Once msLM is pre-trained with Reid and Artetxe’s objective functions [11], we continue with vanilla FT with the NMT objective function.
We use the four FT strategies shown in Figure 1, and we formally define them as follows: Let be the set of all domains where there is parallel data available for model training, for a considered language pair. is a domain for which there is no training data, but only test data is available. In the single-domain FT (vanilla FT) baseline (Figure 1(a)), the model is only fine-tuned with data from one domain . In the in-domain case, both fine-tuning and testing is done using data from domain . In the out-domain case, fine-tuning uses data from , but testing is done on domain . In multi-domain FT (Figure 1(b)), the model is fine-tuned with data from multiple domains . In the in-domain case, this includes data from the target domain as well. In the out-domain case, this trained model is tested on the unseen target domain . In single-domain ITTL (Figure 1(c)), msLM is first fine-tuned with data of domain . This is the intermediate task, which is also called the first stage. Next, this fine-tuned model is further fine-tuned with data from domain . This is the second and last stage222The model can always be fine-tuned in more stages if there is data from different domains. However, we do not consider this option due to lack of computational resources.. In the in-domain case, testing will be done on the same domain . In out-domain case, testing will be done on the unseen domain . In multi-domain ITTL (Figure 1(d)), msLM is first fine-tuned with data from multiple domains in . This model is further fine-tuned with , which is the target domain. In the in-domain case, data of is included in the intermediate stage, as well as in the final stage. Similar to previous setups, in the out-domain case, testing will be done on the unseen domain
4 Experimental Setting
Our experiment setup is inspired from our previous work [10] and we follow the same choices for msLM, languages, datasets, etc.
msLM: We used mBART, as previous research showed it outperformed mT5 [10]. Translation-specific models such as NLLB [21] were omitted, because they cannot be further pre-trained.
Languages: We selected the LRL gu, kn, si, ta along with en and hi (see Table 1). kn is not included in mBART and is at a disadvantage compared to other languages [22]. To fine-tune mBART for kn, we used related language fine-tuning strategy [6] with Telugu.
| Lang. | Family | Script | Joshi | mBART coverage |
| class | in Tokens (M) | |||
| en | Indo European | Latin | 5 | 55608 |
| hi | Indo Aryan | Devanagari | 4 | 1715 |
| gu | Indo Aryan | Gujarati | 1 | 140 |
| kn | Dravidian | Kannada | 1 | – |
| si | Indo Aryan | Sinhala | 1 | 243 |
| ta | Dravidian | Tamil | 3 | 595 |
Datasets: We used both open-domain (CC) and domain-specific (Bible, Gvt, PMI) corpora (see Table 3). Each language has parallel data from three domains. The rather small FLORES dataset is used for out-domain testing only. Bible is the only domain-specific corpus with data for all our languages. The PMI and Gvt corpora are mutually exclusive in our experiments. Therefore, when describing results, we use PMI/Gvt to denote that we use one of these corpora for the considered experiment. Unlike PMI/Gvt and Bible, CC is larger, even for LRLs. However, it is a web-mined corpus, and is considered noisy [24, 25]. For multi-domain experiments, we randomly shuffled the training data from different domains.
Evaluation Metrics: We used SentencePiece BLEU (spBLEU) that was used in FLORES-101 evaluation benchmark [26].
Domain Divergence: We used Jenson-Shanon Divergence (JSD) to quantify domain divergence. Kashyap et al. [27] showed that JSD is a very reliable measure to analyze the performance of a model in a new domain. JSD values between our parallel corpora are given in Table 2.
| Dataset | Gvt test | FLORES test | Bib test | PMI test |
| Gvt train | 0.10 | 0.40 | 0.58 | - |
| CC train | 0.40 | 0.41 | 0.52 | 0.44 |
| Bib train | 0.56 | 0.47 | 0.10 | 0.53 |
| PMI train | - | 0.33 | 0.52 | 0.15 |
Test set specifications are in Table 3. Further details on the data sets, model training, evaluation metrics (including spBLEU signature), and JSD calculation are given in A. We carried out experiments across techniques, languages, and domains for all En-XX pairs. XX-En experiments were carried out for si and kn only to keep the number of experiments at a manageable level.
| Dataset | Domain | Languages | Train Size | Test Size |
| Flores-101 | Open | hi, gu, ka, ta | - | 1k |
| Floresv1 | Open | si | - | 1k |
| CCAligned (CC) | Open | all | 100k | 1k |
| Government (Gvt) | Administrative | si, ta | 50k | 1k |
| PMIndia (PMI) | News | hi | 50k | 1k |
| gu, ka | 25k | 1k | ||
| Web-scraped Bible | Religious | all | 25k | 1k |
5 Results and Discussion
5.1 Pre-training with Bi-text Denoising
Three languages were selected for experiments, considering their representation in mBART (Table 1): kn-not included in mBART, si- 243M tokens and hi- 1757M tokens. We tested for data sizes 1k, 10k, and 25k from individual domains and report the average results. Table 4333see Appendix B for raw results shows that neither of the pre-training techniques was able to surpass the results of vanilla FT on mBART, except in two cases. Even in those two cases, continually pre-trained mBART shows only marginal gains over original mBART.
Interestingly, additional pre-training performed worst for Kannada, which is unseen during the original pretraining of mBART. These results imply that bitext denoising is not a suitable strategy for utilizing auxiliary domain parallel data when the data size is small. Reid and Artetxe [11] reported results might improve if additional monolingual data is incorporated in large quantities; however, this is outside the scope of our research and we leave this for future research. Since pre-training only with parallel data did not yield satisfactory results for vanilla FT, we did not carry out further experiments combining pre-training and ITTL.
5.2 Impact of Data Set Size and Domain Divergence on FT and ITTL
In these experiments, we varied the data set size of both the intermediate and final stages. The smallest size is 1k (termed “small”) and the maximum size is 25k (termed “large”)444These terms for dataset sizes are simply relative. In Ranathunga et al.’s [28], even 25k parallel data is considered an extremely small data setup. However, the maximum data size we can use is bound by domain data availability., where 25k is the largest common size across all three domains (see Table 3). Note that we reused the experiments results of single ITTL of our previous research [10]. For clarity, we used a selected set of EN-XX experiment results for our discussion; full results for EN-XX and XX-EN are in Appendix C.
To identify the best-performing technique under different data setups, we analyzed the results with respect to dataset sizes of the intermediate (IM) and final (FI) stages in the following four combinations (recall that Adelani et al. [8] considered only the large-small data setup): 1) small-small (IM data 1k - TD data 1k); 2) large-small (IM data 25k - TD data 1k); 3) small-large (IM data 1k - TD data 25k); 4) large-large (IM data 25k - TD data 25k).
5.2.1 In-Domain Test Case
In-domain results are reported in Tables 5-8. In this setup, we consider either PMI/Gvt or Bible as the target domain. Baseline results are reported by training mBART with the target domain, and testing with the same. In ITTF experiments, the second stage of fine-tuning is done using this target domain data. In each table, the best technique appears in bold, the second best appears in italic. Worst is grayed out.
| Language | out domain | in domain | ||
| bitex | bitex+mono | bitex | bitex+mono | |
| ka | -0.2 | -0.6 | -0.8 | -0.6 |
| si | -0.2 | -0.4 | -0.5 | 0.3 |
| hi | -0.2 | -0.5 | 0.3 | -0.4 |
| Language | Intermediate task - PMI/Gvt Final task - Bible | Intermediate task - Bible Final task - PMI/Gvt | ||||||
| Single ITTL | Multi FT | Multi ITTL | Baseline | Single ITTL | Multi FT | Multi ITTL | Baseline | |
| si | 17.2 | 16.7 | 16.7 | 17.0 | 20.9 | 21.4 | 22.3 | 20.2 |
| ta | 11.0 | 14.5 | 15.0 | 9.3 | 19.9 | 18.9 | 20.1 | 18.6 |
| gu | 13.0 | 13.3 | 13.2 | 12.9 | 18.8 | 19.4 | 20.1 | 20.9 |
| hi | 15.5 | 15.0 | 15.7 | 14.0 | 18.3 | 18.5 | 19.3 | 18.2 |
| ka | 9.4 | 9.1 | 8.9 | 8.2 | 5.9 | 6.2 | 6.6 | 6.5 |
| Avg | 13.2 | 13.7 | 13.9 | 12.3 | 16.8 | 16.9 | 17.7 | 16.9 |
0cm Language Intermediate task - PMI/Gvt Final task - Bible Intermediate task - Bible Final task - PMI/Gvt Intermediate task - CC Final task - PMI/Gvt Intermediate task - CC Final task - Bible Single ITTL Multi FT Multi ITTL Baseline Single ITTL Multi FT Multi ITTL Baseline Single ITTL Multi FT Multi ITTL Baseline Single ITTL Multi FT Multi ITTL Baseline si 18.9 19.4 19.4 17.0 22.1 19.8 24.5 20.2 24.1 23.7 25.8 20.2 19.3 17.4 18.8 17.0 ta 13.4 13.5 14.5 9.3 18.3 18.3 19.7 18.6 20.0 20.0 21.0 18.6 14.5 14.9 16.4 9.3 gu 14.3 14.6 15.1 12.9 19.9 20.5 21.5 20.9 23.8 24.3 24.6 20.9 13.9 13.3 15.4 12.9 hi 15.3 15.9 16.7 14.0 18.7 18.1 19.2 18.2 20.0 20.8 21.1 18.2 16.1 16.2 16.3 14.0 ka 12.7 12.8 13.8 8.2 12.9 12.5 15.5 6.5 12.9 13.8 14.8 6.5 9.6 9.6 10.0 8.2 Avg 14.9 15.2 15.9 12.3 18.4 17.8 20.1 16.9 20.2 20.5 21.5 16.9 14.7 14.3 15.4 12.3
| Language | Intermediate task - PMI/Gvt Final task - Bible | Intermediate task - Bible Final task - PMI/Gvt | ||||||
| Single ITTL | Multi FT | Multi ITTL | Baseline | Single ITTL | Multi FT | Multi ITTL | Baseline | |
| si | 37.6 | 37.7 | 38.2 | 37.9 | 45.0 | 43.8 | 45.3 | 44.7 |
| ta | 30.5 | 30.9 | 30.7 | 31.0 | 37.1 | 37.7 | 38.1 | 36.8 |
| gu | 27.3 | 27.8 | 27.7 | 27.8 | 38.0 | 37.8 | 37.8 | 37.9 |
| hi | 31.5 | 31.7 | 31.2 | 31.4 | 35.1 | 35.3 | 34.9 | 35.2 |
| ka | 27.0 | 27.4 | 27.6 | 27.8 | 33.8 | 33.8 | 34 | 33.6 |
| Avg | 30.8 | 31.1 | 31.1 | 31.2 | 37.8 | 37.7 | 38.0 | 37.6 |
0cm Language Intermediate task - PMI/Gvt Final task - Bible Intermediate task - Bible Final task - PMI/Gvt Intermediate task - CC Final task - PMI/Gvt Intermediate task - CC Final task - Bible Single ITTL Multi FT Multi ITTL Baseline Single ITTL Multi FT Multi ITTL Baseline Single ITTL Multi FT Multi ITTL Baseline Single ITTL Multi FT Multi ITTL Baseline si 38.1 37.9 37.9 37.9 44.2 43.7 44.0 44.7 44.4 43.9 44.7 44.7 37.9 38.0 37.9 37.9 ta 30.9 30.4 30.7 31.0 37.2 37.0 38.2 36.8 37.2 37.7 37.9 36.8 30.8 30.6 30.7 31.0 gu 27.4 27.7 28.2 27.8 37.6 37.7 37.6 37.9 38.3 37.5 37.6 37.9 26.9 27.8 27.9 27.8 hi 31.6 31.4 31.4 31.4 34.4 35.1 35.1 35.2 35.5 35.4 35.3 35.2 31.5 31.4 31.1 31.4 ka 27.5 26.2 27.2 27.8 32.8 32.7 34.4 33.6 34.3 33.5 33.7 33.6 27.3 26.5 27.4 27.8 Avg 31.1 30.7 31.1 31.2 37.2 37.2 37.9 37.6 37.9 37.6 37.8 37.6 30.9 30.9 31.0 31.2
small-small
According to Table 5, multi-domain ITTL is the best-performing technique. However, on average, its gains over the second-best technique (multi-domain FT, which uses less compute) is less than 1 spBLEU. Hence, under limited computing resources, multi-domain FT the best option.
large-small
Comparing Table 6 against Table 5 (small-small setup), increasing the intermediate task size from 1k to 25k is consistently beneficial. In this data setup, multi-domain ITTL performs better than the other two techniques (more than 1 average spBLEU gains except in one experiment) irrespective of the domain differences, thus confirming Adelani et al.’s [8] observations (see Section 2.1).
small-large and large-large
According to Tables 7 and 8, when there are 25k sentences from the target domain, none of the techniques managed to significantly outperform vanilla FT baseline, thus questioning the benefit of auxiliary domain data.
The above observations are supported by Figure 2 as well 555Graphs for all the experiments are in D. Note that the line corresponding to 0k in the graphs is the vanilla FT baseline. When the target (final) task size is less than 10k, all three techniques are effective over the vanilla FT baseline, but to varying degrees. However, as the target task data set size increases, their results converge to baseline. Even using 100k CC data was not able to make gains in performance when the target domain data set is at least 25k. Thus our results emphasize that Adelani et al.’s [8] observations do not always hold, and it is important to consider the size of the datasets.
An important observation is the results for ka, which is unseen in mBART. Adding an intermediate task with 1k auxiliary data only results in a less than 1 spBLEU gain for this language in the small-small data setup. However, when adding 25k auxiliary data at the intermediate stage results in a significant gain (5.9 spBLEU on average) resulting in the largest gain amongst the languages.
5.2.2 Out-domain Test Case
| Language | Intermediate task - PMI/Gvt Final task - Bible | Intermediate task - Bible Final task - PMI/Gvt | ||||||
| Single ITTL | Multi FT | Multi ITTL | Baseline | Single ITTL | Multi FT | Multi ITTL | Baseline | |
| si | 2.4 | 4.4 | 4.3 | 0.9 | 4.9 | 4.4 | 4.5 | 3.9 |
| ta | 2.2 | 3.7 | 3.6 | 0.7 | 3.9 | 3.7 | 3.9 | 2.6 |
| gu | 6.0 | 7.8 | 7.9 | 2.2 | 8.3 | 7.8 | 8.1 | 7.3 |
| hi | 4.8 | 8.0 | 7.6 | 2.6 | 8.4 | 8.0 | 8.2 | 7.1 |
| ka | 0.5 | 1.1 | 1.1 | 0.3 | 1.0 | 1.1 | 1.2 | 1.0 |
| Avg | 3.2 | 5.0 | 4.9 | 1.3 | 5.3 | 5.0 | 5.2 | 4.4 |
| Language | Intermediate task - PMI/Gvt Final task - Bible | Intermediate task - Bible Final task - PMI/Gvt | ||||||
| Single ITTL | Multi FT | Multi ITTL | Baseline | Single ITTL | Multi FT | Multi ITTL | Baseline | |
| si | 1.9 | 4.6 | 2.8 | 1.9 | 11.3 | 11.3 | 11.3 | 11.2 |
| ta | 2.2 | 3.8 | 2.9 | 2.1 | 9.5 | 9.5 | 9.8 | 9.0 |
| gu | 5.2 | 9.4 | 9.0 | 4.4 | 20.2 | 19.8 | 20.0 | 19.8 |
| hi | 3.4 | 6.9 | 6.5 | 3.2 | 16.9 | 17.4 | 17.8 | 17.3 |
| ka | 2.3 | 3.9 | 3.2 | 2.2 | 14.0 | 14.5 | 14.5 | 14.1 |
| Avg | 3.0 | 5.7 | 4.9 | 2.8 | 14.4 | 14.5 | 14.7 | 14.3 |
0mm Language Intermediate task - PMI/Gvt Final task - Bible Intermediate task - Bible Final task - PMI/Gvt Intermediate task - CC Final task - PMI/Gvt Intermediate task - CC Final task - Bible Single ITTL Multi FT Multi ITTL Baseline Single ITTL Multi FT Multi ITTL Baseline Single ITTL Multi FT Multi ITTL Baseline Single ITTL Multi FT Multi ITTL Baseline si 7.9 11.3 11.3 0.9 5.8 4.2 5.6 3.9 8.9 9.7 10.6 3.9 3.5 8.8 8.4 0.9 ta 5.4 9.4 8.7 0.7 4.9 3.8 4.3 2.6 10.1 12.2 12.4 2.6 7.2 10.8 10.6 0.7 gu 15.3 19.8 19.7 2.2 10.5 9.3 10.2 7.3 16.8 16.1 16.5 7.3 9.6 12.2 11.9 2.2 hi 12.8 17.1 17.1 2.6 8.8 7.1 8.2 7.1 13.3 16.3 16.0 7.1 7.1 14.9 14.9 2.6 ka 7.6 14.3 13.7 0.3 4.5 4.4 5.1 1.0 4.5 4.7 5.6 1.0 0.9 1.7 1.9 0.3 Avg 9.8 14.4 14.1 1.34 6.9 5.8 6.7 4.4 10.7 11.8 12.2 4.4 5.7 9.7 9.5 1.3
| Language | Intermediate task - PMI/Gvt Final task - Bible | Intermediate task - Bible Final task - PMI/Gvt | Intermediate task - CC Final task - PMI/Gvt | Intermediate task - CC Final task - Bible | ||||||||||||
| Single ITTL | Multi FT | Multi ITTL | Baseline | Single ITTL | Multi FT | Multi ITTL | Baseline | Single ITTL | Multi FT | Multi ITTL | Baseline | Single ITTL | Multi FT | Multi ITTL | Baseline | |
| si | 2.4 | 11.6 | 11.5 | 1.9 | 11.9 | 9.9 | 10.5 | 11.2 | 12.1 | 14.1 | 13.3 | 11.2 | 2.0 | 9.5 | 8.8 | 1.9 |
| ta | 2.8 | 9.9 | 6.6 | 2.1 | 10.0 | 11.5 | 11.8 | 9.0 | 11.1 | 14.8 | 14.7 | 9.0 | 2.9 | 11.0 | 6.1 | 2.1 |
| gu | 8.0 | 20.8 | 20.0 | 4.4 | 20.5 | 20.8 | 20.6 | 19.8 | 21.5 | 23.7 | 24.1 | 19.8 | 5.6 | 13.3 | 14.6 | 4.4 |
| hi | 4.3 | 17.6 | 16.5 | 3.2 | 18.4 | 17.2 | 18.0 | 17.3 | 18.3 | 21.5 | 21.3 | 17.3 | 3.8 | 14.5 | 12.6 | 3.2 |
| ka | 3.2 | 14.3 | 8.6 | 2.2 | 14.2 | 14.3 | 15.5 | 14.1 | 15.0 | 16.6 | 16.2 | 14.1 | 2.4 | 4.2 | 3.1 | 2.2 |
| Avg | 4.1 | 14.8 | 12.6 | 2.8 | 15.0 | 14.7 | 15.2 | 14.2 | 15.6 | 18.1 | 17.9 | 14.3 | 3.3 | 10.5 | 9.0 | 2.8 |
We use the results for the FLORES test set to explain our observations in Tables 9-12. In ITTL experiments, either PMI/Gvt or Bible is used as the final stage of fine-tuning. PMI/Gvt, Bible, and/or CC are used at the intermediate stage666Unlike domain-specific datasets s.a Bible, CC corpus is larger even for LRLs. Thus we do not run 1k setup for CC.. Here, baseline refers to vanilla FT with the dataset used for the final task (e.g. if Bible is used as the final task in the ITTL experiments, baseline refers to vanilla FT with Bible and testing with FLORES).
Unlike in the in-domain test case, in the out-domain test case, there exists a large impact of domain divergence. To begin with, vanilla FT baseline performs significantly better (average gain 11.5 BLEU) when PMI/Gvt is used for fine-tuning, compared to using Bible. This observation can be explained with the domain divergence results in Table 2, where the divergence between FLORES and Bible () is higher than that between FLORES and PMI/Gvt (/).
Similarly, for all four data size setups, the best results from both the ITTL strategies are reported when a domain more closer to FLORES is used in final stage fine-tuning. The best result is reported when both the intermediate dataset and the final dataset are from domains closer to the target domain (e.g. CC and PMI/Gvt).
The impact of domain divergence is visible in Figure 3 as well. For example, in single-domain ITTL in Figure 3(b), results for FLORES decrease to baseline when Bible dataset size increases in the final stage. In contrast, in Figure 3(a), when PMI/Gvt is used in the final stage, results for FLORES mostly increase when PMI/Gvt size increases.
For all four data size setups, all three techniques are better than the vanilla FT baseline, but in varying degrees. In order to further explain this observation, we calculated the correlation () between JSD (of the final stage dataset and test set) and the corresponding spBLEU score. We see that the average is for the baseline, while the same is , 0 and for single-domain ITTL, multi-domain ITTL and multi-domain FT (respectively). This indicates that, compared to the vanilla FT baseline, domain divergence is less correlated to the performance when auxiliary domain data is used.
Overall, both multi-domain ITTL and multi-domain FT are better than single-domain ITTL, except in two cases. However, the results difference between multi-domain ITTL and multi-domain FT is marginal. In fact, in more than half of the experiments reported in Tables 9-12, multi-domain ITTL underperforms multi-domain FT. Our summary results also show that the variance of performance for multi-domain FT and multi-domain ITTL is and (respectively). The higher variance of multi-domain ITTL indicates that the final stage fine-tuning in the multi-domain ITTL model is more susceptible to domain divergence compared to multi-domain FT.
6 Ablation Study
6.1 Impact of Up-sampling
In both small-large and large-small dataset setups, one dataset size is significantly smaller than the other. In order to dilute the impact of size difference, we up-sampled777random oversampling (https://t.ly/MLrhd), we use the term up-sampling to mean oversampling the smaller data set to match with the larger dataset. For this experiment, we used CC (the largest dataset) and Bible, and experimented with three languages. As shown in Table 13, we do not see any gains from upsampling.
| Method | CC Size | Bible Size | hi | ka | si | ||||||
| FLORES | Bible | PMI | FLORES | Bible | PMI | FLORES | Bible | Gvt | |||
| Multi FT | 25k | 1k | -6.8 | 0.6 | -4.4 | -1.3 | -0.6 | -0.9 | -4.2 | 0.2 | -7.5 |
| 10k | -1.8 | 0.3 | -1.8 | -0.8 | 0.4 | -0.5 | -1.5 | 0.2 | -2.9 | ||
| 100k | 1k | -13.1 | 1.2 | -9.2 | -3.8 | 3.9 | -2.5 | -8.5 | 1.3 | -13.6 | |
| 10k | -7.9 | 0.6 | -5.2 | -0.7 | 6.0 | -0.6 | -4.8 | 1.1 | -7.9 | ||
| 25k | -4.9 | 0.3 | -3.4 | -1.4 | 2.7 | -0.7 | -2.7 | 1.4 | -2.6 | ||
| Multi ITTL | 25k | 1k | -7.1 | 0.5 | -4.7 | -1.5 | -0.4 | -0.9 | -3.9 | -0.9 | -7.2 |
| 10k | -1.6 | 0.2 | -1.6 | -0.4 | 0.3 | -0.1 | -1.7 | -0.1 | -3.2 | ||
| 100k | 1k | -14 | 0.6 | -9.5 | -4.8 | -0.1 | -2.9 | -9.1 | 0.1 | -13.8 | |
| 10k | -8.1 | 0.6 | -5.7 | -0.6 | 0.2 | 0.0 | -2.0 | -0.6 | -1.7 | ||
| 25k | 2.1 | -0.7 | 1.6 | 0.2 | 1.2 | -0.2 | 0.5 | 0.9 | 0.8 | ||
6.2 Mixing more than two domains
All our previous experiments involved only two domains. In order to investigate the impact of combining multiple domains, we mixed data from CC, Bible, and PMI/Gvt separately for si and ka. As seen in Tables 14 and 15, the inclusion of an extra domain during training does not consistently improve performance due to domain divergence.
| Multi-Domains | Domain Size | ka | si | ||||
| FLORES | Bible | PMI | FLORES | Bible | Gvt | ||
| CC+Bible | 25k+25k | 4.2 | 26.5 | 2.1 | 9.5 | 38.0 | 16.6 |
| CC+PMI/Gvt | 25k+25k | 16.6 | 1.1 | 33.5 | 14.1 | 2.7 | 43.9 |
| CC+PMI/Gvt+Bible | 25k+25k+25k | 14.5 | 25.1 | 31.9 | 13.6 | 37.2 | 43.5 |
| Intermediate Task | Final Stage | ka | si | ||||||
| Multi-Domains | Domain Size | Domain | Size | FLORES | Bible | PMI | FLORES | Bible | Gvt |
| CC+Bible | 25k+25k | Bible | 25k | 3.1 | 27.4 | 1.4 | 8.8 | 37.9 | 15.6 |
| CC+PMI/Gvt+Bible | 25k+25k+25k | 7.0 | 27.8 | 14.4 | 9.1 | 37.6 | 34.3 | ||
| Improvement | 3.9 | 0.4 | 13.0 | 0.3 | -0.3 | 18.7 | |||
| CC+PMI/Gvt | 25k+25k | PMI/Gvt | 25k | 16.2 | 1.8 | 33.7 | 13.3 | 2.3 | 44.7 |
| CC+PMI/Gvt+Bible | 25k+25k+25k | 16.6 | 19.8 | 34.3 | 13.8 | 31.0 | 45.0 | ||
| Improvement | 0.4 | 18.0 | 0.6 | 0.5 | 28.7 | 0.3 | |||
In multi-domain FT (Table 14), when testing on FLORES, combining Bible with CC+PMI/Gvt yielded inferior results. A similar observation holds for the in-domain test case—adding Bible to CC+PMI/Gvt drops the result on PMI/Gvt.
In multi-domain ITTL (Table 15), when testing on FLORES, adding PMI/Gvt to CC+Bible significantly improves the results, but adding Bible to CC+PMI/Gvt yields marginal gains. Similarly, in the in-domain case, adding PMI/Gvt to CC+Bible does not increase the result on Bible, and adding Bible to CC+PMI/Gvt does not increase the result on PMI/Gvt.
7 Conclusions
The optimal way of using auxiliary domain parallel data to build domain-specific LRL-NMT systems on msLMs was an under-explored problem. We presented the first comparative study of two ways to utilize this parallel data—fine-tuning and additional pre-training. Based on our experiments, we are able to provide a set of observations, which will assist NMT researchers in selecting the optimal way to use data from auxiliary domains, without having to run extensive sets of experiments. In the future, we plan to experiment with multilingual FT and ITTL, as well as multilingual and multi-domain FT and ITTL. It would also be worthwhile to explore new pre-training objectives on parallel data as well.
Appendix A Experiment Setup
A.1 Corpus Details (from [10])
PMIndia corpus (PMI)
[29] is a multi-way parallel corpus of 14 languages (13 Indian languages, plus English). The data set contains Indian Prime Minister’s speeches extracted from the Prime Minister of India website888https://www.pmindia.gov.in/en/.
Government corpus (Gvt)
[30] is a multi-way parallel corpus for Sinhala, Tamil, and English. This is a manually curated data set. It consists of data from official government documents (annual reports, committee reports, government institutional websites, and acts of the Parliament).
Bible corpus
We scrape Bible data from web999Sinhala: https://www.wordproject.org/bibles/si/index.htm; and others: https://ebible.org/download.php and then automatically align the sentences (at verse level) using the scripts provided by Nayak et al. [10] This curated multi-way parallel corpus has 25k parallel sentences in kn, gu, hi, ta. Note that Sinhala was scraped from a different website, thus has different content101010 We will be releasing the scripts to create the corpus on acceptance of the paper..
CCAligned (CC)
is an automatically sentence-aligned (bi-text mined) corpus based on the CommonCrawl.
FLORES-101
A.2 Model Training
Our implementation is based on the Fairseq library. All experiments are run with seed 222. We used an Nvidia Volta of 32 GPU RAM.
Training details are as follows: number of epochs =, learning rate = , dropout = , attention dropout = , and a batch size = for training and evaluation.
A.3 Evaluation Metric
spBLEU calculates the BLEU score for the text tokenized using the sentence-piece sub-word model, which has been trained for all the 101 languages in Flores-101 data set (including the languages (except kn) considered in our experiments). Furthermore, Goyal et al. [26] showed that spBLEU functions similarly to BLEU and has a strong correlation with the Chrf++ metric. [32] 121212We use the official implementation provided in the sacreBLEU library111111https://github.com/mjpost/sacreBLEU [33].
The sacreBLEU signature for pre-training with bi-text denoising is:
nrefs:1—case:mixed—eff:no—tok:
spm-flores—smooth:none—version:2.1.0.
The sacreBLEU signature for Single-domain ITTL is:
nrefs:
1—case:mixed—eff:no—tok:spm-flores—smooth:none—
version:2.2.1.
The signature for Multi-domain ITTL is:
nrefs:1—case:mixed—eff:no—tok:flores101—
smooth:none—version:2.3.1.
A.4 Jenson-Shannon Divergence
Jenson-Shannon Divergence (JSD) is calculated between two distributions and using the formula given below:
Here, represents an equally weighted sum of the two distributions (i.e., ). represents the Kullback–Leibler divergence. JSD ranges from 0 to 1 - lower the value, more similar the distributions are.
Divergence is calculated between each pair of train and test sets, for a given language pair. These language-specific values are then averaged to obtain a similarity score independent of the target language.
Following our previous work [10], each corpus was first tokenized (times and numbers were also converted into the tokens <TIME> and <NUMBER>, respectively) using the NLTK package,131313https://www.nltk.org/ and stopwords were removed. The resulting corpora were transformed into a (discrete) frequency distribution over all word tokens. The frequency distributions of each train and test set are then compared using the formula above.
Appendix B Pre-training results
| Pre-training Set | Pre-training | Pre-training Set Size | Fine-tuning Set | Fine-tuning Set Size | ka | si | hi | ||||||
| FLORES | Bible | PMI | FLORES | Bible | Gvt | FLORES | Bible | PMI | |||||
| CC | bitext | 25k | CC | 25k | 1.0 | 0.0 | 0.6 | 8.5 | 1.9 | 14.1 | 12.3 | 5.0 | 11.3 |
| 100k | 100k | 6.2 | 0.1 | 4.5 | 15.4 | 4.1 | 26.4 | 22.7 | 6.8 | 17.8 | |||
| bitext + monolignual | 25k | 25k | 0.5 | 0.0 | 0.4 | 8.4 | 1.8 | 13.3 | 14.0 | 5.4 | 11.9 | ||
| 100k | 100k | 5.8 | 0.2 | 4.1 | 14.6 | 4.1 | 25.9 | 22.9 | 6.9 | 18.4 | |||
| PMI/Gvt | bitext | 1k | PMI/Gvt | 1k | 0.5 | 0.1 | 3.7 | 3.7 | 0.5 | 19.5 | 6.5 | 1.4 | 17.3 |
| 10k | 10k | 9.6 | 1.1 | 28.9 | 8.8 | 0.9 | 39.7 | 14 | 2.2 | 31.7 | |||
| 25k | 25k | 14.3 | 1.9 | 34.3 | 10.8 | 1.1 | 44.9 | 17.2 | 3 | 35.6 | |||
| 50k | 50k | N/A | N/A | N/A | 12.2 | 1.3 | 48.7 | 19.1 | 3.3 | 37.2 | |||
| bitext + monolingual | 1k | 1k | 0.5 | 0.1 | 3.3 | 3.2 | 0.4 | 21.2 | 7.1 | 1.7 | 18.6 | ||
| 10k | 10k | 9.1 | 0.9 | 29.1 | 8.3 | 0.8 | 38.3 | 14 | 2.4 | 31.8 | |||
| 25k | 25k | 12.8 | 1.6 | 33.8 | 10.7 | 1.1 | 44.5 | 17.2 | 3 | 35.6 | |||
| 50k | 50k | N/A | N/A | N/A | 11.7 | 1.3 | 49.2 | 18.7 | 3.3 | 36.8 | |||
| Bible | bitext | 1k | Bible | 1k | 0.3 | 6.7 | 0.2 | 0.7 | 12.6 | 1.1 | 2.4 | 15.9 | 2.1 |
| 10k | 10k | 1.5 | 21.9 | 0.6 | 1.8 | 33.7 | 0.8 | 2.9 | 26.9 | 2.2 | |||
| 25k | 25k | 2.1 | 28 | 0.8 | 1.6 | 38 | 0.7 | 3 | 31.6 | 2.1 | |||
| bitext + monolignual | 1k | 1k | 0.3 | 7.4 | 0.2 | 0.8 | 17.9 | 1.2 | 0.9 | 10.5 | 1.2 | ||
| 10k | 10k | 1.3 | 22.4 | 0.6 | 1.8 | 33.9 | 0.8 | 2.9 | 26.8 | 2 | |||
| 25k | 25k | 1.8 | 28 | 0.7 | 1.7 | 37.9 | 0.7 | 2.7 | 31.4 | 1.9 | |||
Appendix C Tabular Results for Individual Techniques
C.1 Result Tables for EN-XX
| D1 | D1 size | D2 | D2 size | ka | gu | hi | si | ta | ||||||||||
| FLORES | Bible | PMI | FLORES | Bible | PMI | FLORES | Bible | PMI | FLORES | Bible | Gvt | FLORES | Bible | Gvt | ||||
| CC | 25k | PMI/Gvt | 1k | 4.5 | 0.4 | 12.9 | 16.8 | 3.6 | 23.8 | 13.3 | 3.6 | 20 | 8.9 | 1.7 | 24.1 | 10.1 | 5.9 | 20.0 |
| 10k | 12.4 | 1.6 | 30.5 | 19.4 | 4.6 | 34.7 | 16.6 | 3.3 | 31.8 | 10.6 | 1.9 | 39.2 | 9.8 | 3.6 | 33.4 | |||
| 25k | 15.0 | 1.9 | 34.3 | 21.5 | 5.1 | 38.3 | 18.3 | 3.3 | 35.5 | 12.1 | 1.9 | 44.4 | 11.1 | 3.3 | 37.2 | |||
| 50k | N/A | N/A | N/A | N/A | N/A | N/A | 19.9 | 3.8 | 37.0 | 12.8 | 2.0 | 49.7 | 12.1 | 2.8 | 40.1 | |||
| 100k | 1k | 10.2 | 1.1 | 16.2 | 20.9 | 3.7 | 26.0 | 22.1 | 6.1 | 24.0 | 16.3 | 4.6 | 31.0 | 20.3 | 9.4 | 23.9 | ||
| 10k | 14.6 | 1.8 | 30.3 | 22.5 | 5.2 | 35.1 | 20.3 | 4.9 | 32.5 | 15.2 | 4.1 | 41.4 | 15.4 | 6.2 | 34.1 | |||
| 25k | 16.3 | 2.1 | 34.1 | 22.9 | 5.4 | 38.5 | 21.4 | 4.4 | 35.2 | 15.2 | 3.6 | 44.3 | 15.0 | 5.0 | 37.6 | |||
| 50k | N/A | N/A | N/A | N/A | N/A | N/A | 22.5 | 4.4 | 37.0 | 15.2 | 3.5 | 48.7 | 14.9 | 4.5 | 40.5 | |||
| CC | 25k | Bible | 1k | 0.9 | 9.6 | 0.7 | 9.6 | 13.9 | 8.5 | 7.1 | 16.1 | 6.8 | 3.5 | 19.3 | 6.2 | 7.2 | 14.5 | 6.7 |
| 10k | 2.1 | 23.1 | 0.9 | 6.6 | 23.7 | 5.1 | 4.0 | 26.4 | 3.0 | 2.2 | 33.5 | 2.2 | 3.0 | 26.1 | 2.5 | |||
| 25k | 2.4 | 27.3 | 1 | 5.6 | 26.9 | 4.1 | 3.8 | 31.5 | 2.4 | 2.0 | 37.9 | 1.0 | 2.9 | 30.8 | 1.6 | |||
| 100k | 1k | 4.1 | 11.2 | 2.3 | 14.6 | 13.3 | 11.9 | 16.9 | 16.5 | 14.5 | 11.0 | 20.1 | 19.0 | 17.2 | 16.3 | 15.3 | ||
| 10k | 2.8 | 22.9 | 1.4 | 8.3 | 24.2 | 6.9 | 6.6 | 26.8 | 5.5 | 3.9 | 33.4 | 5.4 | 5.9 | 26.2 | 6.0 | |||
| 25k | 2.8 | 27.4 | 1.2 | 6.9 | 27.8 | 5.4 | 4.8 | 31.7 | 3.3 | 3.0 | 37.9 | 2.6 | 4.2 | 31.0 | 3.2 | |||
| PMI/Gvt | 1k | Bible | 1k | 0.5 | 9.4 | 1.0 | 6.0 | 13.0 | 15.6 | 4.8 | 15.5 | 11.4 | 2.4 | 17.2 | 13.1 | 2.2 | 11.0 | 13.7 |
| 10k | 1.7 | 22.0 | 0.9 | 5.1 | 24.2 | 6.6 | 3.3 | 27.2 | 2.9 | 1.9 | 34.1 | 3.2 | 2.3 | 26.3 | 3.7 | |||
| 25k | 2.3 | 27.0 | 1.1 | 5.2 | 27.3 | 5.1 | 3.4 | 31.5 | 2.5 | 1.9 | 37.6 | 1.8 | 2.2 | 30.5 | 2.7 | |||
| 10k | 1k | 4.2 | 12.4 | 12.9 | 11.8 | 14.7 | 26.3 | 10.2 | 15.8 | 24.7 | 5.3 | 17.5 | 29.8 | 3.8 | 13.6 | 23.3 | ||
| 10k | 2.8 | 23.1 | 2.9 | 7.4 | 24.2 | 13.6 | 5.1 | 26.9 | 8.9 | 2.2 | 32.4 | 8.0 | 2.6 | 25.8 | 8.1 | |||
| 25k | 2.9 | 27.1 | 2.1 | 6.5 | 27.7 | 9.1 | 4.2 | 31.5 | 4.6 | 2.1 | 37.8 | 4.4 | 2.5 | 30.6 | 4.8 | |||
| 25k | 1k | 7.6 | 12.7 | 21.3 | 15.3 | 14.3 | 31.8 | 12.8 | 15.3 | 29.3 | 7.9 | 18.9 | 36.7 | 5.4 | 13.4 | 28.4 | ||
| 10k | 3.7 | 23.1 | 6.1 | 9.7 | 23.9 | 20.1 | 6.0 | 27.0 | 12.8 | 3.3 | 33.1 | 14.5 | 3.2 | 24.8 | 13.8 | |||
| 25k | 3.2 | 27.5 | 2.8 | 8.0 | 27.4 | 13.4 | 4.3 | 31.6 | 6.4 | 2.4 | 38.1 | 6.5 | 2.8 | 30.9 | 7.3 | |||
| 50k | 1k | N/A | N/A | N/A | N/A | N/A | N/A | 14.7 | 16.0 | 31.3 | 8.4 | 19.2 | 40.5 | 6.9 | 13.7 | 32.5 | ||
| 10k | N/A | N/A | N/A | N/A | N/A | N/A | 6.7 | 26.7 | 15.8 | 4.3 | 33.5 | 17.8 | 3.8 | 25.0 | 17.3 | |||
| 25k | N/A | N/A | N/A | N/A | N/A | N/A | 4.8 | 31.8 | 8.2 | 3.0 | 37.3 | 8.6 | 3.3 | 30.5 | 9.8 | |||
| Bible | 1k | PMI/Gvt | 1k | 1.0 | 1.6 | 5.9 | 8.3 | 9.6 | 18.8 | 8.4 | 8.6 | 18.3 | 4.9 | 10.3 | 20.9 | 3.9 | 6.1 | 19.9 |
| 10k | 10.0 | 1.7 | 28.5 | 17.3 | 7.5 | 34.5 | 14.7 | 5.3 | 31.4 | 9.3 | 5.5 | 38.6 | 7.0 | 4.6 | 32.7 | |||
| 25k | 14.0 | 2.0 | 33.8 | 20.2 | 6.8 | 38.0 | 16.9 | 4.7 | 35.1 | 11.3 | 3.5 | 45.0 | 9.5 | 3.0 | 37.1 | |||
| 50k | N/A | N/A | N/A | N/A | N/A | N/A | 19.1 | 4.9 | 37.2 | 12.3 | 3.5 | 49.5 | 11.2 | 2.6 | 40.7 | |||
| 10k | 1k | 4.0 | 12.0 | 12.8 | 10.3 | 19.4 | 20.2 | 9.1 | 20.1 | 19.2 | 6.1 | 25.1 | 22.4 | 4.8 | 19.3 | 18.9 | ||
| 10k | 11.4 | 6.0 | 29.7 | 18.1 | 13.9 | 34.8 | 15.4 | 14.2 | 31.4 | 10.0 | 15.1 | 39.3 | 7.9 | 12.2 | 33.2 | |||
| 25k | 14.3 | 4.3 | 34.2 | 20.5 | 12.0 | 37.9 | 17.9 | 10.8 | 34.7 | 11.8 | 12.2 | 44.2 | 9.9 | 10.2 | 37.3 | |||
| 50k | N/A | N/A | N/A | N/A | N/A | N/A | 19.9 | 10.1 | 36.7 | 12.7 | 9.3 | 49.0 | 11.6 | 8.5 | 39.8 | |||
| 25k | 1k | 4.5 | 16.3 | 12.9 | 10.5 | 22.6 | 19.9 | 8.8 | 23.9 | 18.7 | 5.8 | 28.7 | 22.1 | 4.9 | 24.3 | 18.3 | ||
| 10k | 11.6 | 8.4 | 29.8 | 17.8 | 17.4 | 33.7 | 15.4 | 17.2 | 31.2 | 10.0 | 19.6 | 39.1 | 8.0 | 14.9 | 33.6 | |||
| 25k | 14.2 | 6.5 | 32.8 | 20.5 | 14.7 | 37.6 | 18.4 | 15.2 | 34.4 | 11.9 | 14.9 | 44.2 | 10.0 | 13.2 | 37.2 | |||
| 50k | N/A | N/A | N/A | N/A | N/A | N/A | 20 | 12.7 | 36.8 | 12.9 | 12.7 | 48.6 | 11.2 | 11.4 | 39.9 | |||
| D1 + D2 | D1 + D2 size | ka | gu | hi | si | ta | ||||||||||
| FLORES | Bible | PMI | FLORES | Bible | PMI | FLORES | Bible | PMI | FLORES | Bible | Gvt | FLORES | Bible | Gvt | ||
| CC + PMI/Gvt | 25k + 1k | 4.7 | 0.0 | 13.8 | 16.1 | 2.1 | 24.3 | 16.3 | 5.6 | 20.8 | 9.7 | 1.9 | 23.7 | 12.2 | 6.6 | 20.0 |
| 25k + 10k | 13.6 | 0.4 | 30.0 | 22.2 | 3.4 | 34.4 | 19.4 | 5.9 | 31.2 | 12.2 | 2.4 | 37.9 | 13.1 | 7.0 | 32.7 | |
| 25k + 25k | 16.6 | 1.1 | 33.5 | 23.7 | 5.1 | 37.5 | 21.5 | 6.4 | 35.4 | 14.1 | 2.7 | 43.9 | 14.8 | 7.5 | 37.7 | |
| 25k + 50k | N/A | N/A | N/A | N/A | N/A | N/A | 22.9 | 6.6 | 36.9 | 14.9 | 2.8 | 47.9 | 15.3 | 7.4 | 40.0 | |
| 100k + 1k | 6.1 | 0.2 | 11.1 | 15.8 | 1.6 | 24.9 | 23.9 | 7.2 | 23.6 | 15.5 | 4.5 | 30.4 | 18.7 | 9.3 | 21.4 | |
| 100k + 10k | 12.7 | 0.7 | 25.1 | 22.4 | 3.7 | 33.4 | 26.1 | 7.4 | 31.8 | 16.7 | 4.9 | 38.7 | 19.0 | 9.8 | 31.6 | |
| 100k + 25k | 15.0 | 1.0 | 29.8 | 23.5 | 5.0 | 36.8 | 26.9 | 7.5 | 35.2 | 17.3 | 4.6 | 43.2 | 19.6 | 9.4 | 35.3 | |
| 100k + 50k | N/A | N/A | N/A | N/A | N/A | N/A | 27.1 | 7.8 | 36.3 | 17.6 | 4.4 | 45.7 | 19.9 | 8.8 | 37.5 | |
| CC + Bible | 25k + 1k | 1.7 | 9.6 | 1.2 | 12.2 | 13.3 | 8.2 | 14.9 | 16.2 | 12.2 | 8.8 | 17.4 | 14.8 | 10.8 | 14.9 | 8.9 |
| 25k + 10k | 3.6 | 22.4 | 2.0 | 12.8 | 23.9 | 10.7 | 14.8 | 27.0 | 12.6 | 9.8 | 33.1 | 16.9 | 11.3 | 26.1 | 9.4 | |
| 25k + 25k | 4.2 | 26.5 | 2.1 | 13.3 | 27.8 | 11.5 | 14.5 | 31.4 | 11.9 | 9.5 | 38 | 16.6 | 11.0 | 30.6 | 9.5 | |
| 100k + 1k | 4.3 | 7.3 | 2.9 | 12.1 | 13.1 | 11.2 | 23.9 | 16.4 | 19.0 | 15.4 | 18.8 | 26.8 | 18.7 | 16.0 | 17.0 | |
| 100k + 10k | 4.1 | 16.7 | 2.8 | 12.9 | 22.2 | 12.6 | 24.1 | 27.0 | 19.1 | 15.4 | 31.9 | 27.4 | 18.8 | 24.5 | 17.1 | |
| 100k + 25k | 5.8 | 22.6 | 3.5 | 11.9 | 25.0 | 10.2 | 23.8 | 30.6 | 19.3 | 15.4 | 35.7 | 26.8 | 18.1 | 28.3 | 16.2 | |
| PMI/Gvt + Bible (1st) | 1k + 1k | 1.1 | 9.1 | 6.2 | 7.7 | 13.1 | 19.4 | 8.0 | 15.0 | 18.5 | 4.4 | 18.5 | 21.4 | 3.7 | 14.5 | 18.9 |
| 1k + 10k | 3.7 | 22.7 | 11.8 | 8.8 | 23.8 | 20.3 | 7.5 | 26.7 | 18.1 | 4.9 | 33.4 | 22.4 | 3.9 | 25.1 | 17.4 | |
| 1k + 25k | 4.4 | 27.8 | 12.5 | 9.3 | 27.6 | 20.5 | 7.1 | 30.8 | 18.1 | 4.2 | 36.6 | 19.8 | 3.8 | 30.9 | 18.3 | |
| 10k + 1k | 11.2 | 12.6 | 29.3 | 17.2 | 14.2 | 34.0 | 15.0 | 16.7 | 31.0 | 9.1 | 19.4 | 39.3 | 7.5 | 14.9 | 33.7 | |
| 10k + 10k | 11.9 | 22.8 | 29.3 | 17.9 | 24.2 | 34.1 | 14.7 | 27.2 | 30.5 | 9.5 | 33.8 | 39.0 | 7.9 | 25.9 | 33.0 | |
| 10k + 25k | 11.8 | 27.3 | 29.3 | 18.2 | 27.8 | 34.2 | 14.5 | 31.1 | 30.8 | 9.4 | 37.5 | 38.5 | 7.6 | 30.4 | 33.0 | |
| 25k + 1k | 14.5 | 13.0 | 33.8 | 19.8 | 15.0 | 37.8 | 17.4 | 16.5 | 35.3 | 11.3 | 19.2 | 43.8 | 9.5 | 14.7 | 37.7 | |
| 25k + 10k | 14.9 | 23.0 | 33.7 | 20.9 | 24.1 | 37.4 | 17.4 | 27.2 | 35.6 | 11.9 | 33.8 | 44.9 | 9.9 | 25.9 | 37.3 | |
| 25k + 25k | 14.3 | 26.2 | 32.7 | 20.8 | 27.7 | 37.7 | 17.2 | 31.4 | 35.1 | 11.5 | 37.7 | 43.7 | 9.9 | 30.4 | 37.0 | |
| 50k + 1k | N/A | N/A | N/A | N/A | N/A | N/A | 19.4 | 17.2 | 37.4 | 12.4 | 20.0 | 48.6 | 11.2 | 14.1 | 40.1 | |
| 50k + 10k | N/A | N/A | N/A | N/A | N/A | N/A | 20.1 | 26.8 | 37.1 | 12.7 | 33.4 | 48.2 | 11.3 | 25.6 | 39.7 | |
| 50k + 25k | N/A | N/A | N/A | N/A | N/A | N/A | 19.2 | 31.0 | 36.2 | 12.6 | 37.4 | 46.9 | 11.2 | 29.5 | 39.4 | |
| PMI/Gvt + Bible (2nd) | 1k + 1k | 1.1 | 9.1 | 6.2 | 7.9 | 13.3 | 19.2 | 8.0 | 15.0 | 18.5 | 4.3 | 16.7 | 19.5 | 3.7 | 14.5 | 18.9 |
| 1k + 10k | 3.2 | 22.2 | 10.8 | 9.2 | 24.2 | 21.0 | 7.7 | 27.0 | 18.5 | 4.8 | 33.7 | 22.7 | 3.6 | 24.6 | 16.9 | |
| 1k + 25k | 3.9 | 27.4 | 11.3 | 9.4 | 27.8 | 21.2 | 6.9 | 31.7 | 18.2 | 4.6 | 37.7 | 22.3 | 3.8 | 30.9 | 18.3 | |
| 10k + 1k | 10.1 | 11.6 | 28.7 | 16.7 | 14.1 | 33.3 | 15.0 | 16.7 | 31.0 | 8.8 | 18.1 | 37.2 | 7.3 | 13.9 | 32.1 | |
| 10k + 10k | 11.5 | 22.8 | 29.1 | 17.9 | 24.2 | 34.1 | 14.7 | 27.2 | 31.0 | 9.1 | 32.4 | 37.8 | 7.7 | 25.4 | 32.3 | |
| 10k + 25k | 11.7 | 27.0 | 28.9 | 18.2 | 27.8 | 34.2 | 14.6 | 31.6 | 31.2 | 9.3 | 38.2 | 39.3 | 7.6 | 30.7 | 33.1 | |
| 25k + 1k | 14.3 | 12.8 | 33.5 | 19.8 | 14.6 | 37.6 | 17.1 | 15.9 | 34.8 | 11.3 | 19.4 | 43.4 | 9.4 | 13.5 | 36.1 | |
| 25k + 10k | 14.8 | 23.1 | 33.6 | 20.9 | 24.4 | 37.3 | 17.4 | 27.2 | 35.6 | 11.8 | 33.2 | 43.8 | 9.9 | 25.9 | 37.3 | |
| 25k + 25k | 14.3 | 26.2 | 32.7 | 20.8 | 27.7 | 37.7 | 17.6 | 31.4 | 35.1 | 11.6 | 37.9 | 44.3 | 9.9 | 30.4 | 37.0 | |
| 50k + 1k | N/A | N/A | N/A | N/A | N/A | N/A | 18.8 | 15.6 | 36.7 | 12.4 | 18.1 | 45.8 | 11.1 | 13.7 | 39.7 | |
| 50k + 10k | N/A | N/A | N/A | N/A | N/A | N/A | 19.8 | 26.7 | 36.8 | 12.7 | 33.4 | 48.2 | 11.3 | 25.8 | 40.0 | |
| 50k + 25k | N/A | N/A | N/A | N/A | N/A | N/A | 19.2 | 30.5 | 36.9 | 12.6 | 37.4 | 46.9 | 11.2 | 29.5 | 39.4 | |
| D1 + D2 | D1 + D2 size | D2 | D2 size | ka | gu | hi | si | ta | ||||||||||
| FLORES | Bible | PMI | FLORES | Bible | PMI | FLORES | Bible | PMI | FLORES | Bible | Gvt | FLORES | Bible | Gvt | ||||
| CC + PMI/Gvt | 25k + 1k | PMI/Gvt | 1k | 5.6 | 0.1 | 14.8 | 16.5 | 2.2 | 24.6 | 16.0 | 5.6 | 21.1 | 10.6 | 2.2 | 25.8 | 12.4 | 6.7 | 21 |
| 25k + 10k | 10k | 13.7 | 0.4 | 30.4 | 21.7 | 4.7 | 34.7 | 19.3 | 5.9 | 31.2 | 12.7 | 2.8 | 39.2 | 13.4 | 7.1 | 33.1 | ||
| 25k + 25k | 25k | 16.2 | 1.8 | 33.7 | 24.1 | 5.3 | 37.6 | 21.3 | 6.2 | 35.3 | 13.3 | 2.3 | 44.7 | 14.7 | 7.4 | 37.9 | ||
| 25k + 50k | 50k | N/A | N/A | N/A | N/A | N/A | N/A | 22.9 | 6.6 | 36.7 | 14.4 | 2.7 | 49.3 | 13.6 | 6.1 | 40.7 | ||
| 100k + 1k | 1k | 10.0 | 0.6 | 17.8 | 21.9 | 4.0 | 26.8 | 24.1 | 7.0 | 25.7 | 16.9 | 4.8 | 32.6 | 21.1 | 9.9 | 24.2 | ||
| 100k + 10k | 10k | 15.0 | 2.1 | 30.4 | 24.7 | 5.6 | 35.4 | 25.2 | 7.3 | 32.6 | 17.7 | 5.4 | 40.0 | 18.3 | 8.5 | 34.3 | ||
| 100k + 25k | 25k | 17.1 | 2.2 | 34.1 | 24.9 | 5.9 | 38.2 | 25.8 | 6.9 | 35.4 | 17.5 | 5.0 | 45.2 | 17.0 | 7.1 | 38.1 | ||
| 100k + 50k | 50k | N/A | N/A | N/A | N/A | N/A | N/A | 25.2 | 6.4 | 37.2 | 16.9 | 4.7 | 48.5 | 17.0 | 6.5 | 41.1 | ||
| CC + Bible | 25k + 1k | Bible | 1k | 1.9 | 10.0 | 1.2 | 11.9 | 15.4 | 10.0 | 14.9 | 16.3 | 12.3 | 8.4 | 18.8 | 14.5 | 10.6 | 16.4 | 8.8 |
| 25k + 10k | 10k | 3.1 | 22.7 | 1.6 | 14.6 | 24.6 | 11.9 | 13.9 | 27.3 | 11.8 | 9.3 | 33.7 | 16.1 | 10.2 | 26.4 | 9.1 | ||
| 25k + 25k | 25k | 3.1 | 27.4 | 1.4 | 14.6 | 27.9 | 12.6 | 12.6 | 31.1 | 10.7 | 8.8 | 37.9 | 15.6 | 6.1 | 30.7 | 6.7 | ||
| 100k + 1k | 1k | 5.4 | 11.8 | 3.3 | 15.3 | 14.7 | 15.1 | 23.8 | 17.3 | 18.7 | 15.6 | 20.4 | 26.3 | 19.3 | 17.4 | 17.1 | ||
| 100k + 10k | 10k | 3.7 | 22.9 | 1.7 | 17.3 | 23.9 | 14.8 | 24.0 | 27.2 | 18.9 | 11.3 | 34.2 | 18.7 | 12.9 | 26.3 | 12.7 | ||
| 100k + 25k | 25k | 3.6 | 27.1 | 1.7 | 11.0 | 27.9 | 9.1 | 14.1 | 31.6 | 12.2 | 9.1 | 37.4 | 15.7 | 9.8 | 30.1 | 9.8 | ||
| PMI/Gvt + Bible | 1k + 1k | Bible | 1k | 1.1 | 8.9 | 6.1 | 7.9 | 13.2 | 19.4 | 7.6 | 15.7 | 18.2 | 4.3 | 16.7 | 19.4 | 3.6 | 15.0 | 19.1 |
| PMI/Gvt | 1k | 1.2 | 8.6 | 6.6 | 8.1 | 13.0 | 20.1 | 8.2 | 14.9 | 19.3 | 4.5 | 18.3 | 22.3 | 3.9 | 12.0 | 20.1 | ||
| 1k + 10k | Bible | 10k | 2.6 | 22.6 | 7.6 | 9.5 | 24.1 | 21.0 | 6.9 | 27.2 | 18.1 | 4.3 | 34.1 | 21.4 | 3.0 | 25.7 | 13.3 | |
| PMI/Gvt | 1k | 4.0 | 21.7 | 13.7 | 8.9 | 23.8 | 20.3 | 8.3 | 26.3 | 19.5 | 5.2 | 33.2 | 23.1 | 4.4 | 25.0 | 20.0 | ||
| 1k + 25k | Bible | 25k | 3.2 | 27.6 | 5.6 | 9.0 | 27.7 | 19.7 | 6.9 | 31.2 | 17.2 | 2.8 | 38.2 | 14.1 | 2.9 | 30.7 | 12.6 | |
| PMI/Gvt | 1k | 5.1 | 25.6 | 15.5 | 10.2 | 27.2 | 21.5 | 8.2 | 30.1 | 19.2 | 5.6 | 35.6 | 24.5 | 4.3 | 30.2 | 19.7 | ||
| 10k + 1k | Bible | 1k | 9.3 | 12.7 | 26.8 | 16.7 | 14.5 | 33.5 | 14.8 | 16.6 | 31.0 | 8.4 | 19.5 | 37.0 | 7.2 | 14.1 | 32.1 | |
| PMI/Gvt | 10k | 11.4 | 12.2 | 29.3 | 16.9 | 13.2 | 34.1 | 15.0 | 16.3 | 31.6 | 9.2 | 20.1 | 39.1 | 7.3 | 14.8 | 33.9 | ||
| 10k + 10k | Bible | 10k | 9.6 | 23.4 | 25.9 | 17.0 | 24.3 | 33.7 | 14.3 | 27.5 | 30.5 | 6.6 | 33.2 | 31.7 | 7.0 | 25.8 | 31.6 | |
| PMI/Gvt | 10k | 12.0 | 22.4 | 29.4 | 17.7 | 23.1 | 34.2 | 15.4 | 27.2 | 31.5 | 9.6 | 33.8 | 39.1 | 7.7 | 25.6 | 33.2 | ||
| 10k + 25k | Bible | 25k | 7.1 | 27.6 | 17.0 | 17.1 | 27.9 | 33.1 | 13.5 | 31.2 | 30.2 | 9.1 | 37.9 | 39.0 | 5.0 | 30.6 | 25.8 | |
| PMI/Gvt | 10k | 12.2 | 26.0 | 29.5 | 18.5 | 27.6 | 34.6 | 14.9 | 31.1 | 31.8 | 9.7 | 37.1 | 39.7 | 8.0 | 29.9 | 33.6 | ||
| 25k + 1k | Bible | 1k | 13.7 | 13.8 | 32.7 | 19.7 | 15.1 | 37.4 | 17.1 | 16.7 | 34.8 | 11.3 | 19.4 | 43.4 | 8.7 | 14.5 | 36.1 | |
| PMI/Gvt | 25k | 14.5 | 12.9 | 34.0 | 20.0 | 14.0 | 37.8 | 17.8 | 15.4 | 34.9 | 11.3 | 15.4 | 45.3 | 9.8 | 14.7 | 38.1 | ||
| 25k + 10k | Bible | 10k | 13.9 | 23.1 | 32.3 | 20.2 | 24.7 | 36.5 | 17.2 | 27.5 | 35.0 | 11.0 | 33.8 | 43.2 | 8.8 | 26.3 | 36.5 | |
| PMI/Gvt | 25k | 14.7 | 20.8 | 33.7 | 21.2 | 23.1 | 37.8 | 17.7 | 26.8 | 35.4 | 11.7 | 32.8 | 44.8 | 9.9 | 25.7 | 37.3 | ||
| 25k + 25k | Bible | 25k | 8.6 | 27.2 | 20.0 | 20.0 | 28.2 | 36.5 | 16.5 | 31.4 | 34.3 | 11.5 | 37.9 | 44.4 | 6.6 | 30.7 | 29.5 | |
| PMI/Gvt | 25k | 15.5 | 21.4 | 34.4 | 20.6 | 25.0 | 37.6 | 18.0 | 29.5 | 35.1 | 11.8 | 35.7 | 44.0 | 10.5 | 25.3 | 38.2 | ||
| 50k + 1k | Bible | 1k | N/A | N/A | N/A | N/A | N/A | N/A | 18.0 | 17.6 | 36.5 | 11.9 | 20.1 | 45.5 | 10.1 | 14.8 | 39.5 | |
| PMI/Gvt | 50k | N/A | N/A | N/A | N/A | N/A | N/A | 19.9 | 16.8 | 36.9 | 12.3 | 16.5 | 48.6 | 10.7 | 11.4 | 40.8 | ||
| 50k + 10k | Bible | 10k | N/A | N/A | N/A | N/A | N/A | N/A | 15.7 | 27.5 | 32.6 | 12.2 | 34.3 | 47.4 | 10.3 | 26.0 | 39.2 | |
| PMI/Gvt | 50k | N/A | N/A | N/A | N/A | N/A | N/A | 20.0 | 26.7 | 37.0 | 12.8 | 26.5 | 48.9 | 11.5 | 20.8 | 40.7 | ||
| 50k + 25k | Bible | 25k | N/A | N/A | N/A | N/A | N/A | N/A | 18.0 | 30.8 | 35.9 | 8.1 | 38.0 | 35.9 | 7.1 | 30.5 | 31.4 | |
| PMI/Gvt | 50k | N/A | N/A | N/A | N/A | N/A | N/A | 19.8 | 26.6 | 37.3 | 12.8 | 29.9 | 48.5 | 11.6 | 23.9 | 40.5 | ||
| D1 | D1 size | D2 | D2 size | ka | gu | hi | si | ta | ||||||||||
| FLORES | Bible | PMI | FLORES | Bible | PMI | FLORES | Bible | PMI | FLORES | Bible | Gvt | FLORES | Bible | Gvt | ||||
| CC | 25k | PMI/Gvt | 1k | 0.2 | -0.4 | 0.9 | -0.7 | -1.5 | 0.5 | 3.0 | 2.0 | 0.8 | 0.8 | 0.2 | -0.4 | 2.1 | 0.7 | 0.0 |
| 25k | 10k | 1.2 | -1.2 | -0.5 | 2.8 | -1.2 | -0.3 | 2.8 | 2.6 | -0.6 | 1.6 | 0.5 | -1.3 | 3.3 | 3.4 | -0.7 | ||
| 25k | 25k | 1.6 | -0.8 | -0.8 | 2.2 | 0.0 | -0.8 | 3.2 | 3.1 | -0.1 | 2.0 | 0.8 | -0.5 | 3.7 | 4.2 | 0.5 | ||
| 25k | 50k | N/A | N/A | N/A | N/A | N/A | N/A | 3.2 | 2.8 | -0.1 | 2.1 | 0.8 | -1.8 | 3.2 | 4.6 | -0.1 | ||
| 100k | 1k | -4.1 | -0.9 | -5.1 | -5.1 | -2.1 | -1.1 | 3.0 | 1.1 | -0.4 | -0.8 | -0.1 | -0.6 | -1.6 | -0.1 | -2.5 | ||
| 100k | 10k | -1.9 | -1.1 | -5.2 | -0.1 | -1.5 | -1.7 | 1.8 | 2.5 | -0.7 | 1.5 | 0.8 | -2.7 | 3.6 | 3.6 | -2.5 | ||
| 100k | 25k | -1.3 | -1.1 | -4.3 | 0.6 | -0.4 | -1.7 | 5.8 | 3.1 | 0.0 | 2.1 | 1.0 | -1.1 | 4.6 | 4.4 | -2.3 | ||
| 100k | 50k | N/A | N/A | N/A | N/A | N/A | N/A | 4.6 | 3.4 | -0.7 | 2.4 | 0.9 | -3.0 | 5.0 | 4.3 | -3.0 | ||
| CC | 25k | Bible | 1k | 0/8 | 0.0 | 0.5 | 2.6 | -0.6 | -0.3 | 7.8 | 0.1 | 5.4 | 5.3 | -1.9 | 8.6 | 3.6 | 0.4 | 2.2 |
| 25k | 10k | 1.5 | -0.7 | 1.1 | 6.2 | 0.2 | 5.6 | 10.8 | 0.6 | 9.6 | 7.6 | -0.4 | 14.7 | 8.3 | 0.0 | 6.9 | ||
| 25k | 25k | 1.8 | -0.8 | 1.1 | 7.7 | 0.9 | 7.4 | 10.7 | -0.1 | 9.5 | 7.5 | 0.1 | 15.6 | 8.1 | -0.2 | 7.9 | ||
| 100k | 1k | 0.2 | -3.9 | 0.6 | -2.5 | -0.2 | -0.7 | 7.0 | -0.1 | 4.5 | 4.4 | -1.3 | 7.8 | 1.5 | -0.3 | 1.7 | ||
| 100k | 10k | 1.3 | -6.2 | 1.4 | 4.6 | -2.0 | 5.7 | 17.5 | 0.2 | 13.6 | 11.5 | -1.5 | 22.0 | 12.9 | -1.7 | 11.1 | ||
| 100k | 25k | 3.0 | -4.8 | 2.3 | 5.0 | -2.8 | 4.8 | 19.0 | -1.1 | 16.0 | 12.4 | -2.2 | 24.2 | 13.9 | -2.7 | 13.0 | ||
| PMI/Gvt | 1k | Bible | 1k | 0.6 | -0.3 | 5.2 | 1.7 | 0.1 | 3.8 | 3.2 | -0.5 | 7.1 | 2.0 | 1.3 | 8.3 | 1.5 | 3.5 | 5.2 |
| 1k | 10k | 2.0 | 0.7 | 10.9 | 3.7 | -0.4 | 13.7 | 4.2 | -0.5 | 15.2 | 3.0 | -0.7 | 19.2 | 1.6 | -1.2 | 13.7 | ||
| 1k | 25k | 2.1 | 0.8 | 11.4 | 4.1 | 0.3 | 15.4 | 3.7 | -0.7 | 15.6 | 2.3 | -1.0 | 18.0 | 1.6 | 0.4 | 15.6 | ||
| 10k | 1k | 7.0 | 0.2 | 16.4 | 5.4 | -0.5 | 7.7 | 4.8 | 0.9 | 6.3 | 3.8 | 1.9 | 9.5 | 3.7 | 1.3 | 10.4 | ||
| 10k | 10k | 9.1 | -0.3 | 26.4 | 10.5 | 0.0 | 20.5 | 9.6 | 0.3 | 21.6 | 7.3 | 1.4 | 31.0 | 5.3 | 0.1 | 24.9 | ||
| 10k | 25k | 8.9 | 0.2 | 27.2 | 11.7 | 0.1 | 25.1 | 10.3 | -0.4 | 26.2 | 7.3 | -0.3 | 34.1 | 5.1 | -0.2 | 28.2 | ||
| 25k | 1k | 6.9 | 0.3 | 12.5 | 4.5 | 0.7 | 6.0 | 4.6 | 1.2 | 6.0 | 3.4 | 0.3 | 7.1 | 4.1 | 1.3 | 9.3 | ||
| 25k | 10k | 11.2 | -0.1 | 27.6 | 11.2 | 0.2 | 17.3 | 11.4 | 0.2 | 22.8 | 8.6 | 0.7 | 30.4 | 6.7 | 1.1 | 23.5 | ||
| 25k | 25k | 11.1 | -1.3 | 29.9 | 12.8 | 0.3 | 24.3 | 12.9 | -0.2 | 28.7 | 9.1 | -0.4 | 37.2 | 7.1 | -0.5 | 29.7 | ||
| 50k | 1k | N/A | N/A | N/A | N/A | N/A | N/A | 4.7 | 1.2 | 6.1 | 4.0 | 0.8 | 8.1 | 4.3 | 0.4 | 7.6 | ||
| 50k | 10k | N/A | N/A | N/A | N/A | N/A | N/A | 13.4 | 0.1 | 21.3 | 8.4 | -0.1 | 30.4 | 7.5 | 0.6 | 22.4 | ||
| 50k | 25k | N/A | N/A | N/A | N/A | N/A | N/A | 14.4 | -0.8 | 28.0 | 9.6 | 0.1 | 38.3 | 7.9 | -1.0 | 29.6 | ||
| Bible | 1k | PMI/Gvt | 1k | 0.1 | 7.5 | 0.3 | -0.4 | 3.7 | 0.4 | -0.4 | 6.4 | 0.2 | -0.6 | 6.4 | -1.4 | -0.2 | 8.4 | -1.0 |
| 1k | 10k | 0.1 | 9.9 | 0.2 | -0.6 | 6.6 | -1.2 | 0.3 | 11.4 | -0.4 | -0.5 | 12.6 | -1.4 | 0.3 | 9.3 | -0.6 | ||
| 1k | 25k | 0.3 | 10.8 | -0.3 | -0.4 | 7.8 | -0.4 | 0.2 | 11.2 | -0.3 | 0.0 | 15.9 | -1.6 | -0.1 | 10.5 | -1.0 | ||
| 1k | 50k | N/A | N/A | N/A | N/A | N/A | N/A | -0.3 | 10.7 | -0.5 | 0.1 | 14.6 | -3.7 | -0.1 | 11.1 | -1.0 | ||
| 10k | 1k | -0.8 | 10.2 | -2 | -1.1 | 4.8 | 0.8 | -1.4 | 6.9 | -0.7 | -1.3 | 8.6 | 0.3 | -1.2 | 5.3 | -2.0 | ||
| 10k | 10k | 0.1 | 16.8 | -0.6 | -0.2 | 10.3 | -0.7 | -0.7 | 13.0 | -0.4 | -0.9 | 17.3 | -1.5 | -0.2 | 13.2 | -0.9 | ||
| 10k | 25k | 0.5 | 18.8 | -0.6 | 0.4 | 12.4 | -0.6 | -0.5 | 16.4 | 0.9 | 0.0 | 21.0 | -0.4 | 0.0 | 15.7 | 0.0 | ||
| 10k | 50k | N/A | N/A | N/A | N/A | N/A | N/A | -0.1 | 16.6 | 0.1 | 0.0 | 24.1 | -0.8 | -0.3 | 17.3 | 0.2 | ||
| 25k | 1k | -0.6 | 11.1 | -1.6 | -1.1 | 5.2 | 1.3 | -1.9 | 7.8 | -0.5 | -1.2 | 9.0 | 0.2 | -1.1 | 6.6 | 0.0 | ||
| 25k | 10k | 0.1 | 18.6 | -0.9 | -2.3 | 15.8 | -3.7 | -0.8 | 14.4 | 0.0 | -0.7 | 18.6 | 0.2 | -0.4 | 15.8 | -0.5 | ||
| 25k | 25k | 0.1 | 19.7 | -0.1 | 0.3 | 13.0 | 0.1 | -0.8 | 16.2 | 0.7 | -0.3 | 23.0 | 0.1 | -0.1 | 17.2 | -0.2 | ||
| 25k | 50k | N/A | N/A | N/A | N/A | N/A | N/A | -0.8 | 17.8 | 0.1 | -0.3 | 24.7 | -1.7 | 0.0 | 18.1 | -0.5 | ||
| Stage 1 | Stage 2 | ka | gu | hi | si | ta | ||||||||||||||
| D1 | D1 size | D2 | D2 size | D | D size | FLORES | Bible | PMI | FLORES | Bible | PMI | FLORES | Bible | PMI | FLORES | Bible | Gvt | FLORES | Bible | Gvt |
| CC | 25k | PMI/Gvt | 1k | PMI/Gvt | 1k | 0.9 | 0.1 | 1.0 | 0.4 | 0.1 | 0.3 | 0.9 | 0.3 | 2.1 | 0.2 | 0.1 | 1.0 | -0.3 | 0.0 | 0.3 |
| 25k | 10k | 10k | 0.1 | 0.0 | 0.4 | -0.5 | 1.3 | 0.3 | 0.5 | 0.4 | 1.3 | 0.3 | 0.1 | 0.4 | -0.1 | 0.0 | 0.0 | |||
| 25k | 25k | 25k | -0.4 | 0.7 | 0.2 | 0.4 | 0.2 | 0.1 | -0.8 | -0.4 | 0.8 | -0.1 | -0.1 | 0.2 | -0.2 | -0.2 | -0.1 | |||
| 25k | 50k | 50k | N/A | N/A | N/A | N/A | N/A | N/A | -0.5 | -0.1 | 1.4 | -1.7 | -1.3 | 0.7 | 0.0 | 0.0 | 0.0 | |||
| 100k | 1k | 1k | 3.9 | 0.4 | 6.7 | 6.1 | 2.4 | 1.9 | 1.4 | 0.3 | 2.2 | 2.4 | 0.6 | 2.8 | 0.2 | -0.2 | 2.1 | |||
| 100k | 10k | 10k | 2.3 | 1.4 | 5.3 | 2.3 | 1.9 | 2.0 | 1.0 | 0.5 | 1.3 | -0.7 | -1.3 | 2.7 | -0.9 | -0.1 | 0.8 | |||
| 100k | 25k | 25k | 2.1 | 1.2 | 4.3 | 1.4 | 0.9 | 1.4 | 0.2 | 0.4 | 2.0 | -2.6 | -2.3 | 2.8 | -1.1 | -0.6 | 0.2 | |||
| 100k | 50k | 50k | N/A | N/A | N/A | N/A | N/A | N/A | -0.7 | 0.3 | 2.8 | -2.9 | -2.3 | 3.6 | -1.9 | -1.4 | 0.9 | |||
| CC | 25k | Bible | 1k | Bible | 1k | 0.2 | 0.4 | 0.0 | -0.3 | 2.1 | 1.8 | -0.4 | 1.4 | -0.3 | -0.2 | 1.5 | -0.1 | 0.0 | 0.1 | 0.1 |
| 25k | 10k | 10k | -0.5 | 0.3 | -0.4 | 1.8 | 0.7 | 1.2 | -0.5 | 0.6 | -0.8 | -1.1 | 0.3 | -0.3 | -0.9 | 0.3 | -0.8 | |||
| 25k | 25k | 25k | -1.1 | 0.9 | -0.7 | 1.3 | 0.1 | 1.1 | -0.7 | -0.1 | -1.0 | -4.9 | 0.1 | -2.8 | -1.9 | -0.3 | -1.2 | |||
| 100k | 1k | 1k | 1.1 | 4.5 | 0.4 | 3.2 | 1.6 | 3.9 | 0.2 | 1.6 | -0.5 | 0.6 | 1.4 | 0.1 | -0.1 | 0.9 | -0.3 | |||
| 100k | 10k | 10k | -0.4 | 6.2 | -1.1 | 4.4 | 1.7 | 2.2 | -4.1 | 2.3 | -8.7 | -5.9 | 1.8 | -4.4 | -0.1 | 0.2 | -0.2 | |||
| 100k | 25k | 25k | -2.2 | 4.5 | -1.8 | -0.9 | 2.9 | -1.1 | -6.3 | 1.7 | -11.1 | -8.3 | 1.8 | -6.4 | -9.7 | 1.0 | -7.1 | |||
| PMI/Gvt | 1k | Bible | 1k | Bible | 1k | 0.0 | -0.2 | -0.1 | 0.0 | -0.1 | 0.2 | 0.0 | 0.0 | -0.1 | -0.1 | 0.5 | 0.2 | -0.4 | 0.7 | -0.3 |
| PMI/Gvt | 1k | 0.1 | -0.5 | 0.4 | 0.4 | -0.1 | 0.7 | 0.1 | -0.2 | 0.9 | 0.2 | -2.5 | 1.2 | 0.2 | -0.1 | 0.8 | ||||
| 1k | 10k | Bible | 10k | -0.6 | 0.4 | -3.2 | 0.3 | -0.1 | 0.0 | -0.5 | 0.4 | -1.3 | -0.6 | 1.1 | -3.6 | -0.8 | 0.2 | -0.4 | ||
| PMI/Gvt | 1k | 0.3 | -1.0 | 1.9 | 0.1 | 0.0 | 0.0 | 0.3 | -0.2 | 0.7 | 0.5 | -0.1 | 2.6 | 0.8 | -0.4 | 1.4 | ||||
| 1k | 25k | Bible | 25k | -0.7 | 0.2 | -5.7 | -0.4 | -0.1 | -1.5 | -1.8 | 0.5 | -8.2 | -0.9 | -0.2 | -5.7 | -0.4 | -0.5 | -1.0 | ||
| PMI/Gvt | 1k | 0.7 | -2.2 | 3.0 | 0.9 | -0.4 | 1.0 | 1.4 | -1.0 | 4.7 | 0.5 | -0.7 | 1.4 | 1.1 | -0.7 | 1.1 | ||||
| 10k | 1k | Bible | 1k | -0.8 | 1.1 | -1.9 | 0.0 | 0.4 | 0.2 | -0.4 | 1.4 | -0.2 | -0.1 | 0.2 | 0.0 | -0.2 | -0.1 | 0.0 | ||
| PMI/Gvt | 10k | 0.2 | -0.4 | 0.0 | -0.3 | -1.0 | 0.1 | 0.1 | 0.7 | -0.2 | -0.2 | -0.1 | 0.2 | 0.0 | -0.4 | 0.6 | ||||
| 10k | 10k | Bible | 10k | -1.9 | 0.6 | -3.2 | -0.9 | 0.1 | -0.4 | -2.5 | 0.8 | -6.1 | -0.7 | 0.4 | -0.7 | -0.4 | 0.3 | -0.5 | ||
| PMI/Gvt | 10k | 0.1 | -0.4 | 0.1 | -0.2 | -1.1 | 0.1 | 0.1 | 0.0 | 0.1 | -0.2 | -0.3 | 0.2 | 0.7 | 0.0 | 1.0 | ||||
| 10k | 25k | Bible | 25k | -4.6 | 0.6 | -11.9 | -1.1 | 0.1 | -1.1 | -0.2 | -0.3 | -0.3 | -2.6 | -0.1 | -7.3 | -1.1 | -0.4 | -1.0 | ||
| PMI/Gvt | 10k | 0.4 | -1.3 | 0.2 | 0.3 | -0.2 | 0.4 | 0.3 | -0.4 | 1.2 | 0.4 | -0.5 | 0.6 | 0.4 | 0.0 | 1.0 | ||||
| 25k | 1k | Bible | 1k | -0.6 | 1.0 | -0.8 | -0.1 | 0.5 | -0.2 | 0.0 | 0.0 | 0.0 | -0.7 | 1.0 | 0.0 | 0.0 | 0.8 | 0.0 | ||
| PMI/Gvt | 25k | 0.0 | -0.1 | 0.2 | 0.2 | -1.0 | 0.0 | 0.0 | -3.8 | 1.5 | 0.3 | 0.0 | 0.4 | 0.4 | -1.1 | -0.4 | ||||
| 25k | 10k | Bible | 10k | -0.9 | 0.0 | -1.3 | -0.7 | 0.3 | -0.8 | -0.8 | 0.6 | -0.6 | -1.1 | 0.4 | -0.8 | -0.2 | 0.3 | -0.6 | ||
| PMI/Gvt | 25k | -0.2 | -2.2 | 0.0 | 0.3 | -1.0 | 0.4 | -0.2 | -1.0 | -0.1 | 0.0 | -0.2 | 0.0 | 0.3 | -0.4 | -0.2 | ||||
| 25k | 25k | Bible | 25k | -5.7 | 1.0 | -12.7 | -0.8 | 0.5 | -1.2 | -0.1 | 0.0 | 0.1 | -3.3 | 0.3 | -7.5 | -1.1 | 0.0 | -0.8 | ||
| PMI/Gvt | 25k | 1.2 | -4.8 | 1.7 | -0.2 | -2.7 | -0.1 | 0.3 | -2.0 | 0.3 | 0.6 | -5.1 | 1.2 | 0.8 | -1.9 | 0.0 | ||||
| 50k | 1k | Bible | 1k | N/A | N/A | N/A | N/A | N/A | N/A | -0.5 | 2.0 | -0.3 | -1.0 | 1.1 | -0.2 | -0.8 | 2.0 | -0.2 | ||
| PMI/Gvt | 50k | N/A | N/A | N/A | N/A | N/A | N/A | -0.1 | -3.5 | 0.0 | -0.5 | -2.7 | 0.7 | 0.5 | -0.4 | -0.5 | ||||
| 50k | 10k | Bible | 10k | N/A | N/A | N/A | N/A | N/A | N/A | -0.5 | 0.9 | -0.8 | -1.0 | 0.2 | -0.8 | -4.1 | 0.8 | -4.2 | ||
| PMI/Gvt | 50k | N/A | N/A | N/A | N/A | N/A | N/A | 0.1 | -6.9 | 0.7 | 0.2 | -4.8 | 1.0 | -0.1 | -0.1 | -0.1 | ||||
| 50k | 25k | Bible | 25k | N/A | N/A | N/A | N/A | N/A | N/A | -4.5 | 0.6 | -11.0 | -4.1 | 1.0 | -0.8 | -1.2 | 0.3 | -1.0 | ||
| PMI/Gvt | 50k | N/A | N/A | N/A | N/A | N/A | N/A | 0.2 | -7.5 | 1.6 | 0.4 | -5.6 | 1.1 | 0.6 | -4.4 | 1.1 | ||||
| D1 | D1 size | D2 | D2 size | ka | gu | hi | si | ta | ||||||||||
| F | B | P | F | B | P | F | B | P | F | B | G | F | B | G | ||||
| CC | 25k | PMI/Gvt | 1k | 1.1 | -0.3 | 1.9 | -0.3 | -1.4 | 0.8 | 2.7 | 2.0 | 1.1 | 1.7 | 0.5 | 1.7 | 2.3 | 0.8 | 1.0 |
| 25k | 10k | 1.3 | -1.2 | -0.1 | 2.3 | 0.1 | 0.0 | 2.7 | 2.6 | -0.6 | 2.1 | 0.9 | 0.0 | 3.6 | 3.5 | -0.3 | ||
| 25k | 25k | 1.2 | -0.1 | -0.6 | 2.6 | 0.2 | -0.7 | 3.0 | 2.9 | -0.2 | 1.2 | 0.4 | 0.3 | 3.6 | 4.1 | 0.7 | ||
| 25k | 50k | N/A | N/A | N/A | N/A | N/A | N/A | 3.0 | 2.8 | -0.3 | 1.6 | 0.7 | -0.4 | 1.5 | 3.3 | 0.6 | ||
| 100k | 1k | -0.2 | -0.5 | 1.6 | 1.0 | 0.3 | 0.8 | 2.0 | 0.9 | 1.7 | 0.6 | 0.2 | 1.6 | 0.8 | 0.5 | 0.3 | ||
| 100k | 10k | 0.4 | 0.3 | 0.1 | 2.2 | 0.4 | 0.3 | 4.9 | 2.4 | 0.1 | 2.5 | 1.3 | -1.4 | 2.9 | 2.3 | 0.2 | ||
| 100k | 25k | 0.8 | 0.1 | 0.0 | 2.0 | 0.5 | -0.3 | 4.4 | 2.5 | 0.2 | 2.3 | 1.4 | 0.9 | 2.0 | 2.1 | 0.5 | ||
| 100k | 50k | N/A | N/A | N/A | N/A | N/A | N/A | 2.7 | 2.0 | 0.2 | 1.7 | 1.2 | -0.2 | 2.1 | 2.0 | 0.6 | ||
| CC | 25k | Bible | 1k | 1.0 | 0.4 | 0.5 | 2.3 | 1.5 | 1.5 | 7.8 | 0.2 | 5.5 | 4.9 | -0.5 | 8.3 | 3.4 | 1.9 | 2.1 |
| 25k | 10k | 1.0 | -0.4 | 0.7 | 8.0 | 0.9 | 6.8 | 9.9 | 0.9 | 8.8 | 7.1 | 0.2 | 13.9 | 7.2 | 0.3 | 6.6 | ||
| 25k | 25k | 0.7 | 0.1 | 0.4 | 9.0 | 1.0 | 8.5 | 8.8 | -0.4 | 8.3 | 6.8 | 0.0 | 14.6 | 3.2 | -0.1 | 5.1 | ||
| 100k | 1k | 1.3 | 0.6 | 1.0 | 0.7 | 1.4 | 3.2 | 6.9 | 0.8 | 4.2 | 4.6 | 0.3 | 7.3 | 2.1 | 1.1 | 1.8 | ||
| 100k | 10k | 0.9 | 0.0 | 0.3 | 9.0 | -0.3 | 7.9 | 17.4 | 0.4 | 13.4 | 7.4 | 0.8 | 13.3 | 7.0 | 0.1 | 6.7 | ||
| 100k | 25k | 0.8 | -0.3 | 0.5 | 4.1 | 0.1 | 3.7 | 9.3 | -0.1 | 8.9 | 6.1 | -0.5 | 13.1 | 5.6 | -0.9 | 6.6 | ||
| PMI/Gvt | 1k | Bible | 1k | 0.6 | -0.5 | 5.1 | 1.9 | 0.2 | 3.8 | 2.8 | 0.2 | 6.8 | 1.9 | -0.5 | 6.3 | 1.4 | 4.0 | 5.4 |
| 1k | 10k | 0.9 | 0.6 | 6.7 | 4.4 | -0.1 | 14.4 | 3.6 | 0.0 | 15.2 | 2.4 | 0.0 | 18.2 | 0.7 | -0.6 | 9.6 | ||
| 1k | 25k | 0.9 | 0.6 | 4.5 | 3.8 | 0.4 | 14.6 | 3.1 | -0.3 | 14.7 | 0.9 | 0.6 | 12.3 | 0.7 | 0.2 | 9.9 | ||
| 10k | 1k | 5.1 | 0.3 | 13.9 | 4.9 | -0.2 | 7.2 | 4.6 | 0.8 | 6.3 | 3.1 | 2.0 | 7.2 | 3.4 | 0.5 | 8.8 | ||
| 10k | 10k | 6.8 | 0.3 | 23.0 | 9.6 | 0.1 | 20.1 | 9.2 | 0.6 | 21.6 | 4.4 | 0.8 | 23.7 | 4.4 | 0.0 | 23.5 | ||
| 10k | 25k | 4.2 | 0.5 | 14.9 | 10.6 | 0.2 | 24.0 | 9.3 | -0.3 | 25.6 | 7.0 | 0.1 | 34.6 | 2.5 | 0.0 | 21.0 | ||
| 25k | 1k | 6.1 | 1.1 | 11.4 | 4.4 | 0.8 | 5.6 | 4.3 | 1.4 | 5.5 | 3.4 | 0.5 | 6.7 | 3.3 | 1.1 | 7.7 | ||
| 25k | 10k | 10.2 | 0.0 | 26.2 | 10.5 | 0.8 | 16.4 | 11.2 | 0.5 | 22.2 | 7.7 | 0.7 | 28.7 | 5.6 | 1.5 | 22.7 | ||
| 25k | 25k | 5.4 | -0.3 | 17.2 | 12.0 | 0.8 | 23.1 | 12.2 | -0.2 | 27.9 | 9.1 | -0.2 | 37.9 | 3.8 | -0.2 | 22.2 | ||
| 50k | 1k | N/A | N/A | N/A | N/A | N/A | N/A | 3.3 | 1.6 | 5.2 | 3.5 | 0.9 | 5.0 | 3.2 | 1.1 | 7.0 | ||
| 50k | 10k | N/A | N/A | N/A | N/A | N/A | N/A | 9.0 | 0.8 | 16.8 | 7.9 | 0.8 | 29.6 | 6.5 | 1.0 | 21.9 | ||
| 50k | 25k | N/A | N/A | N/A | N/A | N/A | N/A | 13.2 | -1 | 27.7 | 5.1 | 0.7 | 27.3 | 3.8 | 0.0 | 21.6 | ||
| Bible | 1k | PMI/Gvt | 1k | 0.2 | 7.0 | 0.7 | -0.2 | 3.4 | 1.3 | -0.4 | 6.4 | 0.2 | -0.4 | 8.0 | 1.4 | 0.0 | 5.9 | 0.2 |
| 1k | 10k | 1.4 | 10.5 | 0.8 | -0.4 | 5.7 | -0.4 | 0.3 | 11.4 | -0.4 | -0.1 | 14.6 | 0.5 | 0.3 | 10.2 | 1.2 | ||
| 1k | 25k | 0.5 | 10.9 | 0.2 | -0.2 | 7.2 | -0.2 | 0.5 | 11.8 | 0.2 | 0.0 | 11.9 | 0.3 | 0.3 | 11.7 | 1.0 | ||
| 1k | 50k | N/A | N/A | N/A | N/A | N/A | N/A | 0.3 | 12.3 | 0.2 | 0.0 | 13.0 | -0.9 | -0.5 | 8.8 | 0.1 | ||
| 10k | 1k | 0.0 | 9.7 | 0.9 | -1.4 | 4.4 | 0.1 | -1.6 | 6.6 | -1.1 | -0.9 | 8.1 | 0.7 | -0.4 | 5.7 | 1.1 | ||
| 10k | 10k | 0.6 | 16.4 | -0.3 | -0.4 | 9.2 | -0.6 | -0.7 | 13.0 | -0.9 | -0.4 | 18.7 | -0.2 | -0.2 | 13.4 | 0.0 | ||
| 10k | 25k | 0.4 | 16.5 | -0.5 | 0.7 | 11.1 | -0.1 | -0.5 | 16.4 | 0.9 | -0.1 | 20.6 | 0.6 | 0.0 | 15.5 | 0.0 | ||
| 10k | 50k | N/A | N/A | N/A | N/A | N/A | N/A | 0.2 | 16.7 | 0.4 | 0.1 | 17.2 | -0.1 | -0.1 | 12.3 | 0.9 | ||
| 25k | 1k | 0.6 | 9.3 | 2.6 | -0.3 | 4.6 | 1.6 | -1.7 | 6.9 | -0.6 | -0.2 | 6.9 | 2.4 | -0.6 | 5.9 | 1.4 | ||
| 25k | 10k | 0.6 | 17.6 | -0.3 | 0.7 | 10.2 | 0.9 | -0.9 | 13.9 | -0.4 | -0.3 | 17.5 | 0.6 | 0.0 | 15.0 | 0.0 | ||
| 25k | 25k | 1.3 | 14.9 | 1.6 | 0.1 | 10.3 | 0.0 | -1.2 | 16.2 | 0.7 | -0.1 | 20.8 | -0.2 | 0.5 | 12.1 | 1.0 | ||
| 25k | 50k | N/A | N/A | N/A | N/A | N/A | N/A | -0.8 | 18.3 | -0.6 | -0.1 | 17.2 | -0.1 | 0.4 | 12.5 | 0.6 | ||
| Fine-tuning Set | Fine-tuning Set size | ka | si | hi | ||||||
| FLORES | Bible | PMI | FLORES | Bible | Gvt | FLORES | Bible | PMI | ||
| CC | 25k | 0.9 | 0.0 | 0.5 | 8.1 | 1.7 | 13.2 | 12.4 | 4.4 | 11.2 |
| 100k | 4.7 | 0.2 | 3.7 | 14.8 | 4.1 | 26.1 | 21.9 | 6.7 | 17.2 | |
| PMI/Gvt | 1k | 1.0 | 0.1 | 6.5 | 3.9 | 0.5 | 20.2 | 7.1 | 1.5 | 18.2 |
| 10k | 10.4 | 1.3 | 29.4 | 8.8 | 1.0 | 38.2 | 14.7 | 2.6 | 31.4 | |
| 25k | 14.1 | 1.7 | 33.6 | 11.2 | 1.2 | 44.7 | 17.3 | 2.9 | 35.2 | |
| 50k | N/A | N/A | N/A | 12.4 | 1.6 | 49.5 | 19.0 | 3.4 | 37.0 | |
| Bible | 1k | 0.3 | 8.2 | 0.2 | 0.9 | 17.0 | 1.3 | 2.6 | 14.0 | 2.3 |
| 10k | 1.6 | 22.6 | 0.7 | 1.8 | 33.3 | 0.8 | 2.9 | 26.9 | 2.2 | |
| 25k | 2.2 | 27.8 | 0.9 | 1.9 | 37.9 | 0.9 | 3.2 | 31.4 | 2.0 | |
C.2 Result Tables for XX-EN
| D1 | D1 size | D2 | D2 size | ka | si | ||||
| FLORES | Bible | PMI | FLORES | Bible | Gvt | ||||
| CC | 0k | Bible | 1k | 0.0 | 0.8 | 0.0 | 2.4 | 13.8 | 1.8 |
| 10k | 0.9 | 19.0 | 0.6 | 2.5 | 30.3 | 1.4 | |||
| 25k | 1.4 | 28.5 | 1.1 | 2.5 | 35.1 | 1.1 | |||
| 25k | 1k | 0.7 | 2.2 | 0.4 | 6.1 | 15.3 | 9.4 | ||
| 10k | 1.4 | 22.2 | 0.9 | 3.2 | 31.1 | 2.4 | |||
| 25k | 1.6 | 28.5 | 1.2 | 3.0 | 35.4 | 1.4 | |||
| 100k | 1k | 5.1 | 10.7 | 4.4 | 9.4 | 17.0 | 16.9 | ||
| 10k | 2.7 | 23.8 | 2.1 | 4.1 | 31.1 | 4.2 | |||
| 25k | 2.2 | 29.0 | 1.7 | 3.6 | 35.3 | 2.2 | |||
| D1 + D2 | D1 + D2 size | D2 | D2 size | ka | si | ||||
| FLORES | Bible | PMI | FLORES | Bible | Gvt | ||||
| CC + Bible | 25k + 1k | Bible | 1k | 0.5 | 3.1 | 0.5 | 9.7 | 16.3 | 13.5 |
| 25k + 10k | 10k | 2.5 | 22.7 | 1.7 | 8.8 | 30.0 | 13.4 | ||
| 25k + 25k | 25k | 2.8 | 28.2 | 1.9 | 8.1 | 34.9 | 13.0 | ||
| 100k + 1k | 1k | 0.6 | 6.3 | 0.7 | 13.9 | 17.5 | 21.3 | ||
| 100k + 10k | 10k | 1.4 | 22.4 | 1.9 | 14.0 | 30.1 | 21.6 | ||
| 100k + 25k | 25k | 3.4 | 26.1 | 3.4 | 13.4 | 33.3 | 20.7 | ||
| D1 + D2 | D1 + D2 size | D2 | D2 size | ka | si | ||||
| FLORES | Bible | PMI | FLORES | Bible | Gvt | ||||
| CC + Bible | 25k + 1k | Bible | 1k | 0.6 | 3 | 0.5 | 9.7 | 17 | 13.5 |
| 25k + 10k | 10k | 2.7 | 23.5 | 2 | 7.1 | 31.2 | 11.7 | ||
| 25k + 25k | 25k | 2.4 | 28.2 | 1.6 | 5.8 | 35.4 | 8.9 | ||
| 100k + 1k | 1k | 5.4 | 9.8 | 4.8 | 14 | 18.3 | 21.3 | ||
| 100k + 10k | 10k | 4.2 | 24.5 | 3.2 | 11.4 | 30.8 | 20.2 | ||
| 100k + 25k | 25k | 3.2 | 29.4 | 2.5 | 5.5 | 35.7 | 7.3 | ||
Appendix D Graphical Results of Performance
D.1 Single-domain ITTL Performance Plots
D.2 Muti-domain FT Performance Plots
D.3 Multi ITTL Performance Plots